### v1

# Parameter-Efficient Fine-Tuning of Large Language Models for Unit Test Generation: An Empirical Study

Abstract
The advent of large language models (LLMs) like GitHub Copilot has significantly enhanced programmers’ productivity, particularly in code generation. However, these models often struggle with real-world tasks without fine-tuning. As LLMs grow larger and more performant, fine-tuning for specialized tasks becomes increasingly expensive. Parameter-efficient fine-tuning (PEFT) methods, which fine-tune only a subset of model parameters, offer a promising solution by reducing the computational costs of tuning LLMs while maintaining their performance. Existing studies have explored using PEFT and LLMs for various code-related tasks and found that the effectiveness of PEFT techniques is task-dependent. The application of PEFT techniques in unit test generation remains underexplored. The state-of-the-art is limited to using LLMs with full fine-tuning to generate unit tests. This paper investigates both full fine-tuning and various PEFT methods, including LoRA, (IA)3, and prompt tuning, across different model architectures and sizes. We use well-established benchmark datasets to evaluate their effectiveness in unit test generation. Our findings show that PEFT methods can deliver performance comparable to full fine-tuning for unit test generation, making specialized fine-tuning more accessible and cost-effective. Notably, prompt tuning is the most effective in terms of cost and resource utilization, while LoRA approaches the effectiveness of full fine-tuning in several cases.
Index Terms:
large language models, parameter-efficient fine-tuning, test generation, unit tests
I Introduction
Large language models (LLMs) have been developed for various tasks, including general language and code-related activities, significantly enhancing programmers’ performance. Training LLMs on extensive code bases like GitHub give them seemingly magical capabilities. However, despite their proficiency in multiple code-related tasks, LLMs still struggle with cost-effectiveness in their training. The traditional approach to handling these specialized tasks has been fine-tuning the model specifically for each task. Given that LLMs often have billions, if not trillions, of parameters, traditional fine-tuning, which adjusts all parameters, is extremely computationally expensive and resource-intensive. However, state-of-the-art approaches [1] still rely on energy-costly full fine-tuning.
Advancements in instruction-tuned LLMs have encouraged investigations into various prompting and tuning techniques, e.g., [2, 3, 4], that do not require fine-tuning. This challenge has led to the development of multiple techniques for fine-tuning larger models with limited resources, commonly referred to as parameter-efficient fine-tuning (PEFT).
The application of various PEFT methods has been explored for code-related generation tasks, such as code completion, code summarization, and more [5, 6, 7, 8, 9, 10, 11, 12, 13]. Evaluations of the PEFT methods show that they could achieve performance comparable to full fine-tuning while significantly reducing the computational burden by strategically fine-tuning only a chosen subset of parameters during the code-related generation process. However, the effectiveness of PEFT techniques on different tasks varies.
LLMs have been used to generate unit test cases and assertion statements [14, 12, 2, 3, 4, 1]. However, to the best of our knowledge, the area of unit test generation using PEFT remains largely unexplored. Therefore, in this work, we empirically evaluate the performance of various PEFT methods for unit test generation. We focus on answering the following research questions:
  • RQ1: How well does PEFT perform on unit test generation?
  • RQ2: What is the relation between resource utilization and performance of PEFT in unit test generation?
To answer the research questions, we compared full fine-tuning with three popular PEFT methods: LoRA (Low-Rank Adaptation) [15], (IA)3 (Infused Adapter by Inhibiting and Amplifying Inner Activations) [16], and prompt tuning [17]. We conducted these comparisons using ten LLMs from three open-source, decoder-only LLM families. The LLMs have varying architectures and sizes, ranging from 350 million to 16 billion parameters. We measured and compared the unit test generation performance of full fine-tuning and PEFT methods using well-established benchmark datasets rooted in real-world unit test cases.
Our results show that LoRA is generally the most reliable method for improving the quality of the generated unit tests and is, in many cases, on par with or better than full fine-tuning. Conversely, (IA)3 appears to be the least effective, often resulting in minimal improvements while consistently retaining prior knowledge in the LLMs. Prompt tuning demonstrates significant variability, excelling with some models, particularly larger ones, while underperforming with others. This variability highlights the importance of considering model size and characteristics when choosing the appropriate tuning method. Prompt tuning can be considered ”low-hanging fruit,” while LoRA should be employed for greater stability. Our contributions are as follows:
  • We conducted the first empirical study to extensively evaluate training LLMs using various PEFT methods—LoRA, (IA)3, and prompt tuning—for unit test generation across a wide range of LLM models.
  • We offered a comprehensive comparison of PEFT methods versus full fine-tuning.
  • Our findings yield practical guidelines highlighting the potential gains and limitations of the various PEFT methods and full fine-tuning in resource utilization and performance.
The rest of the paper is organized as follows. We provide preliminaries in Section II. Section III introduces related work. We explain our experimental design in Section IV. Section V presents the experimental results. Section VI discusses our results and limitations. Section VII concludes the study and proposes future work.
II Preliminaries
II-A LLM training
The classic language modeling task involves predicting the next word in a sequence, given the preceding words. The model is trained in an autoregressive manner. A typical example is the GPT model [18]. Training an autoregressive model involves fitting a function $f:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$, where $N$ represents the sequence length of tokens, each token having a dimension of $D$. The model aims to estimate the conditional probability distribution, denoted $p$, which can be expressed as:
$p(x_{1:T}\mid c)=\prod_{t=1}^{T}p(x_{t}\mid x_{1:t-1},c)$ (1)
where $x_{t}\in\mathbb{R}^{D}$ represents the $t$’th discrete token observation, $c$ denotes the initial context or input sequence, and $T$ signifies the length of the sequence of predicted tokens [19, Chapter 22].
LLMs are normally pre-trained on a huge amount of data, giving them a broad understanding of language patterns and semantics. This pre-training phase allows LLMs to capture intricate linguistic structures and contextual nuances, empowering them with a robust foundation for various natural language processing tasks.
Fine-tuning is a method in which a pre-trained model can be adapted to a more specific task, such as auto-completing source code. The process involves adjusting the weights of the pre-trained model using task-specific data, thereby enhancing its performance on a target task. When all parameters are tuned, this is often referred to as “full fine-tuning” [20]. If we imagine the model as a web of knowledge, the full fine-tuning loosens all the connections and lets them adapt entirely to the new domain. This flexibility allows for potentially superior performance but comes at a cost:
  • Overfitting: The model can become overly reliant on the training data, potentially struggling with unseen examples [21].
  • Computational cost: Training all parameters requires significant computational resources [20].
  • Catastrophic forgetting: The model might forget some valuable knowledge it acquired during pre-training [22, 23].
II-B PEFT Methods
Parameter-efficient fine-tuning (PEFT) methods provide an alternative way of fine-tuning a pre-trained model by training so-called adapters [24]. Here, most (if not all) parameters of the pre-trained model are frozen, and only a small number of parameters are trained. These adapters are often orders of magnitude smaller than the full model, making sharing, storing, and loading especially convenient. Moreover, some PEFT methods even facilitate the utilization of mixed-task batches, enabling distinct processing for different examples within a batch [17].
PEFT methods can be classified into three main classes of methods: additive, selective, and reparametrization-based [25]. Additive methods can be further broken down into adapters  [^1] The term ”adapter” is also used as the name of the first proposed adapter module for NLP. and soft prompts.
II-B1 Additive
This approach focuses on augmenting the pre-trained model with additional parameters or layers. Only the newly added components are trained, keeping the original model weights frozen. Popular techniques within the additive category are (IA)3 [16], adapters [26], and prompt tuning [17].
Adapters
Adapter, or “bottleneck adapter,” introduced by Houlsby et al. [26] was arguably the first PEFT method. The class of “adapters” introduces small, fully connected networks placed after specific sub-layers within the model. These help the model adapt to the new task without significantly altering the core representation learned during pre-training.
Soft prompts
These techniques focus on creating the so-called “soft prompts.” Prompt tuning, introduced by Lester et al. [17], fine-tuned a subset of the model’s input embeddings, allowing the model to adapt to new tasks without altering the model’s parameters. The prefix tuning expands the idea to all model layers [27].
II-B2 Selective
When using selective approaches, only a specific subset of the pre-trained model’s parameters is fine-tuned. This approach requires carefully selecting the parameters to be updated, ensuring that they are most relevant to the new task. An example is BitFit by Ben Zaken et al. [28], which only updates the bias parameters.
II-B3 Reparametrization-based
This category utilizes techniques that reformulate the original model’s parameters into a more efficient trainable representation. This makes it possible to work with high-dimensional matrices while simultaneously reducing the number of trainable parameters. LoRA [15] is a common reparametrization-based method.
III Related Work
This section presents state-of-the-art studies generating unit tests using LLMs and PEFT approaches to enhance code-related tasks.
III-A Unit test generation using LLMs
Researchers have turned to neural techniques to automatically generate unit tests to align with real-world best practices in unit test generation. Tufano et al. [1] propose an automated unit test case generation tool called AthenaTest. They construct a real-world focal method and developer-written test case dataset named Methods2Test 222 https://github.com/microsoft/methods2test and train a sequence-to-sequence transformer to generate unit test cases.
Advancements in instruction-tuned LLMs have also encouraged several investigations of various prompting techniques, as these do not require any fine-tuning [2, 3, 4]. Yuan et al. [3] propose ChatTester, a ChatGPT-based unit test generation approach, which leverages ChatGPT itself to improve the quality of its generated tests. Schäfer et al. [4] conducted a large-scale empirical evaluation on the effectiveness of LLMs for automated unit test generation without additional training or manual effort. Similarly to ChatTester, they also use ChatGPT (GPT-3.5-Turbo) and propose a re-prompting approach named TestPilot that automatically generates unit tests for all API functions in an npm package. Siddiq et al. [2] use the HumanEval and SF110 datasets to benchmark how well Codex, GPT-3.5-Turbo, and StarCoder can automatically generate unit tests. They evaluate the models based on compilation rates, test correctness, test coverage, and test smells.
III-B PEFT methods on code-related tasks
While prompting methods are gaining traction, they often require significant manual effort to design complex templates and rely on expensive enterprise-grade LLMs like ChatGPT. Consequently, most cutting-edge methods for automatic code generation still involve computationally expensive full fine-tuning. One alternative approach to full fine-tuning is proposed by Steenhoek et al. [29]. They use Reinforcement Learning from Static Quality Metrics (RLSQM). As reinforcement learning is a very expensive fine-tuning method, they also see the value of using PEFT methods like LoRA to facilitate the training.
PEFT has emerged as a powerful technique for enhancing code generation with LLMs. Researchers are investigating how these models perform on code tasks using PEFT techniques as a more efficient alternative to full fine-tuning and PEFT techniques’ effectiveness for different code-related tasks. Results show that the effectiveness of PEFT techniques is task-dependent and not transferrable. For example, Ayupov and Chirkova [5] evaluated adapters and LoRA on various tasks, finding them effective for code understanding but less so for code generation compared to full fine-tuning. Wang et al. [7] reports that prompt tuning of CodeBERT and CodeT5 outperforms full fine-tuning for defect prediction, code summarization, and code translation tasks. Liu et al. [13] evaluated adapter, LoRA, Prefix tuning, and Mix-And-Match adapter (MAM) [30] for defect detection, code clone detection, and code summarization. They concluded that “PEFT may achieve comparable or higher performance than full fine-tuning in code understanding tasks, but they may exhibit slightly weaker performance in code generation tasks.” Similarly, Weyssow et al. [8] evaluate text-to-code generation using LoRA, (IA)3, prompt tuning, and prefix tuning. They compare performance across models of varying sizes (up to 7 billion parameters) against full fine-tuning of smaller models (less than 350 million parameters) and concluded that LoRA is the most effective one.
To our knowledge, there is currently no systematic knowledge of how various PEFT methods behave in the domain of unit test generation. As existing studies show that PEFT techniques’ effectiveness is task-dependent, it is, therefore, necessary to empirically evaluate their effectiveness for unit test generation.
IV Experimental design
The main objective of this study is to gain insight into the performance of fine-tuning LLMs using PEFT for unit test generation.
IV-A Research Questions
To address RQ1 (”How well does PEFT perform on unit test generation?”), we evaluate several different PEFT methods on unit test generation. We also compare the PEFT methods with full fine-tuning. Our study covers a wide range of language models for code of various sizes, spanning 350 million to 16 billion parameters. We assess the quality of the generated test cases produced by different PEFT methods by scoring against real-world human-written reference tests and evaluating the test cases’ syntactic correctness. Further, we examine whether applying PEFT methods for unit test generation adversely affects the pre-trained models’ code generation performance, also known as catastrophic forgetting [22, 23]. This evaluation will help us understand if there are any trade-offs between applying fine-tuning and PEFT methods for unit test generation and maintaining performance on other code-related tasks.
Regarding RQ2 (”What is the relation between resource utilization and performance of PEFT in unit test generation?”), we want to understand PEFT methods’ performance gains in relation to their resource utilization and requirements. Thus, we analyze the efficiency of PEFT methods in terms of the computational resources they use and their impact on the performance outcomes in unit test generation.
IV-B Fine-tuning methods selection
In line with other empirical studies on PEFT [8, 31], we select the popularly used PEFT methods: LoRA [15], (IA)3 [16], and prompt tuning [17]. They are readily available in popular libraries and exemplify most of the primary approaches within PEFT.
(a) Diagram of LoRA (based on [15]).
(b) Diagram of (IA)3 (based on [16]).
(c) Diagram of prompt tuning (based on [17]).
Figure 1: PEFT methods architecture illustration diagrams based on the reference implementation of each introductory paper. Striped modules are frozen. Magenta modules are trained.
  • LoRA: A reparametrization-based method that updates a weight matrix by training a new matrix that is decomposed into a product of two matrices of lower rank: $\delta W=W_{A}\cdot W_{B}$, where $W_{A}\in\mathbb{R}^{\text{in}\times r}$ and $W_{B}\in\mathbb{R}^{r\times\text{out}}$ [15]. As demonstrated in Figure 1(a), the decomposed matrix is then added to the original frozen K and V projection matrices in the attention modules. All other modules are kept frozen.
  • (IA)3: An additive method that focuses on fine-tuning specific intermediate activations within the model [16]. As illustrated in Figure 1(b), this is done by learning three new vectors: $l_{v},l_{k},l_{ff}$; rescaling key, value, and hidden feed-forward activations, respectively.
  • Prompt tuning: A soft prompts method that fine-tunes a subset of the model’s input embeddings. Exemplified in Figure 1(c), the model input embeddings are prepended with a trainable tensor $P\in\mathbb{R}^{N\times D}$ (known as a “soft prompt”) [17]. Each element of $P$ is also referred to as a “virtual token.” During training, these “virtual tokens” are updated while all the rest of the model parameters are kept frozen.
IV-C Models
For our experiments, we select a diverse set of open-source decoder-only model families trained on code to evaluate the performance and efficiency of PEFT methods. The chosen models span a range of sizes and architectures and are trained on diverse datasets and training objectives. This offers extensive insights into the interplay between various model capacities and PEFT methods. Further, all models are ensured to support the Java programming language to facilitate a fair comparison of our benchmarking datasets described in Section IV-D.
  • Salesforce CodeGen Models: To democratize program synthesis, Salesforce AI Research developed CodeGen [32]. Pioneering in multi-turn program synthesis, CodeGen presents a variety of sizes, ranging from 350 million to 16 billion parameters. These models have undergone training across a diverse array of both programming and natural languages. The models were trained on The pile [33], followed by a code subset of Google’s BigQuery datasets Cloud [34]. An updated version CodeGen2 [35] is also available in sizes 1 billion to 16 billion, trained on The Stack dataset [36] using both causal language modeling (CLM) and file-level span corruption. We select the smallest model of the first generation CodeGen family (CodeGen-350M-multi), along with all model sizes of the second generation (1 billion, 3.7 billion, 7 billion, and 16 billion parameters).
  • Meta LLaMA Models: Code Llama [37], derived from Llama 2 [38], is a specialized model further fine-tuned on the code-specific parts of its pre-training data. With enhanced coding capabilities, it can generate code, interpret natural language prompts, assist in code completion, and aid debugging. It comes in three variants. One is optimized for instructions. Another for Python. We select the base variant CodeLlama-7B, as it is designed for general code synthesis and understanding and supports a wide range of popular languages, including Java.
  • BigCode StarCoder Models: This model family includes the 15.5 billion parameters StarCoderBase model [39], along with an updated model generation named StarCoder2 [40], available in 3 billion, 7 billion, and 15 billion parameters. StarCoderBase was trained on 80+ programming languages from The Stack dataset [36]. The updated StarCoder2 is trained on 619 programming languages and other data sources, such as GitHub pull requests, Kaggle notebooks, and code docs. We select all model generations and sizes in this model family.
IV-D Datasets
IV-D1 Methods2Test
To ensure our experiments’ validity and practical applicability, we rely on Methods2Test [1]. It is the largest publicly available corpus of Java unit test cases and corresponding focal methods. This dataset contains 780,000 test cases sourced from 91,000 open-source repositories on GitHub. These real-world test cases represent the complexity and variability inherent in software development, making them an ideal choice for evaluating automated test generation methods. By leveraging Methods2Test, we can assess the performance of the various PEFT methods and models in generating high-quality unit tests that align with real-world coding practices and challenges.
One of the advantages promised by various PEFT methods is the reduction in the amount of data needed for effective training. Thus, we opt to downsize the Methods2Test dataset to enhance efficiency without compromising the representativeness of the data. Specifically, we adopt a strategy wherein only one focal function is selected from each class in the dataset. By doing so, we maintain the essence and diversity of the original dataset while significantly reducing its size. Further, the Methods2Test dataset is distributed in deconstructed components. Hence, we assemble it according to the structure seen in LABEL:lst:data-structure and repackage it as Methods2Testsmall [^3] Available at https://huggingface.co/datasets/andstor/methods2test_small. The resulting dataset contains 7,44K training, 953 validation, and 1.02K testing tests.
// Focal class
public class Calculator {
// Focal method
public float add(float value) {
this.previousValue = this.currentValue;
this.currentValue += value;
return this.currentValue;
// Constructor signatures
public Calculator();
public Calculator(float value);
// Method signatures
public float subtract(float value);
public float multiply(float value);
public float divide(float value);
public void reset();
// Fields
public float currentValue;
public float previousValue;
// Test class
public class CalculatorTest {
// Test case
@Test
public void testAdd() {
Calculator calc = new Calculator();
assertEquals(5.0, calc.add(5.0f), 0.001);
assertEquals(7.0, calc.add(2.0f), 0.001);
} fm+fc+c+m+ft+tcTargetSource
LISTING I: Dataset components: focal method (fm), focal class name (fc), constructor signatures (c), method signatures, fields (f), test case (t), test class name (tc). Data is assembeld according to fm+fc+c+m+f+t+tc. The code is split after the test case signature. Everything above the red line is the source input for unit test generation. The code below the red line is the target output of the unit test generation.
IV-D2 HumanEval-X
HumanEval is a benchmarking dataset for code generation, first introduced by Chen et al. [41]. It consists of 164 hand-written programming problems written in Python. For each programming problem, a function signature, docstring, body, and several unit tests are provided. Zheng et al. [42] extended the dataset to multiple languages, named HumanEval-X. We use the Java subset of the HumanEval-X dataset to evaluate potential adverse fine-tuning effects.
IV-E Training
As the training dataset Methods2Testsmall only includes method signatures for non-focal methods, the unit test generation task can be considered a sequence-to-sequence (seq2seq) problem. For training the decoder-only models described in Section IV-C, we, therefore, model the seq2seq task as an autoregressive problem. This is done by only calculating loss based on the target code, i.e., the unit test code. We minimize the following autoregressive cross-entropy loss function:
$\mathcal{L}=-\sum_{t=1}^{T}\mathds{1}_{t}\cdot\log p(x_{t}\mid x_{1:t-1}),$ (2)
where:
$\mathds{1}_{i}=\begin{cases}1,&\text{if }x_{i}\neq-100\\ 0,&\text{otherwise}.\end{cases}$ (3)
The model is fed the source and target concatenated together (see LABEL:lst:data-structure). The token values for the source are set to -100. The indicator function $\mathds{1}_{i}$ is used to ignore the source in the loss computation.
TABLE I: Model-agnostic hyperparameters for training.
Hyperparameter Method Value
[…]
Optimizer - AdamW
LR schedule - Linear
LR warmup ratio - 0.1
Batch size - 1
Gradient accumulation steps - 8
# Epochs - 3
Precision - Mixed
Learning rate Full fine-tuning 5E-5
LoRA 3E-4
(IA)3 3E-4
Prompt tuning 3E-3
[…]
Alpha LoRA 32
Dropout LoRA 0.1
Rank LoRA 16
Virtual tokens Prompt tuning 20
We use the Transformers [43] and PEFT libraries [44] from Hugging Face to conduct all training. We use NVIDIA A100 GPUs and leverage DeepSpeed ZeRO-Offload [45] where necessary. We use the hyperparameters shown in Table I for training. Like other studies, e.g., [8], we select hyperparameter values that align with those specified in the original papers of the respective PEFT techniques. Weyssow et al. [8] fine-tuned their models for up to five epochs. Our results show that the losses of most of our fine-tuned models are stable after three epochs [^4] Available at https://huggingface.co/andstor/peft-unit-test-generation-experiments. Regarding the unspecified hyperparameters, we leave them to library defaults. Following defaults of the PEFT library, rank decomposition for LoRA is applied to the attention layers. The same attention layers and the default feed-forward modules (see Table II) are targeted for (IA)3. The selected models have context windows ranging from 2048 to 16,384. To make the experiments align with each other, we limit the number of tokens to the smallest window of 2048.
TABLE II: Model-specific hyperparameters for training.
Hyperparameter Method Model Value
Targeted attention modules LoRA, (IA)3 codegen-350M-multi qkv_proj
Salesforce/codegen2-1B_P qkv_proj
Salesforce/codegen2-3_7B_P qkv_proj
Salesforce/codegen2-7B_P qkv_proj
Salesforce/codegen2-16B_P qkv_proj
meta-llama/CodeLlama-7b-hf q_proj, v_proj
bigcode/starcoderbase c_attn
bigcode/starcoder2-3b q_proj, v_proj
bigcode/starcoder2-7b q_proj, v_proj
bigcode/starcoder2-15b q_proj, v_proj
Targeted feedforward modules (IA)3 codegen-350M-multi fc_out
Salesforce/codegen2-1B_P fc_out
Salesforce/codegen2-3_7B_P fc_out
Salesforce/codegen2-7B_P fc_out
Salesforce/codegen2-16B_P fc_out
meta-llama/CodeLlama-7b-hf down_proj
bigcode/starcoderbase mlp.c_proj
bigcode/starcoder2-3b q_proj, c_proj
bigcode/starcoder2-7b q_proj, c_proj
bigcode/starcoder2-15b q_proj, c_proj
IV-F Generation
For generating, we provide the source part of the data (see LABEL:lst:data-structure) as input to the models, while the target part (original code) is used as the ground truth. The generated unit test code is then compared to the ground truth to calculate their similarities. We implement a stopping strategy based on matching braces to generate well-formed functions. For incomplete unit tests, we follow the protocol from [2] where we iteratively delete lines (from bottom to top), add one to two curly brackets, and check if the syntax is valid using an ANTLR v4-based [46] Java parser [47]. We only evaluate samples that are well-formed.
We also apply DeepSpeed-Inference [48] to facilitate model parallel (MP) to fit the larger models that would otherwise not fit in GPU memory. Table III shows the hyperparameter details used for generation.
TABLE III: Hyperparameters for generation.
Hyperparameters Value
Do sample False
Temperature 0
Top p 0
Frequency penalty 0
Max length 2048
IV-G Metrics
IV-G1 Syntactic Correctness
Following Yuan et al. [3], we measure the syntactic correctness of the generated unit tests, ensuring that they conform to the grammar of the Java programming language. After the fixing described in Section IV-F, we employ the same parser as described in Section IV-F. This involves parsing the generated tests and verifying that they are free of syntax errors. We then report the percentage of unit tests with valid syntax for each model.
IV-G2 Similarity
BLEU score [49] is a very common metric used to evaluate natural language machine translation. It measures the similarity of the machine-translated text to the reference translations. However, Ren et al. [50] recognized that it neglects important syntactic and semantic features of programming languages. To remedy this, they developed CodeBLEU [50]. Based on n-gram matching used in BLEU, they further augment the scoring by leveraging abstract syntax trees (AST) and code semantics via data flow. To evaluate the primary performance of the various PEFT methods, we use CodeBLEU to measure the similarity between the generated unit test code and the ground truth.
IV-G3 Coverage
We also execute the generated tests for the HumanEval-Xjava, and calculate code coverage. Specifically, we calculate the pass@1 metric [41] and instruction and branch coverage using the Java Code Coverage Library JaCoCo [51]. Instruction coverage measures the percentage of executed instructions, while branch coverage evaluates the percentage of executed control flow branches in the program.
V Experimental Results
This section presents the detailed results of each research question.
V-A RQ1: Tuning method performance
We present the performance results of the unit test generation for the various models and tuning methods in Table IV.
TABLE IV: Comparison of syntactical validity and CodeBLEU scores from experiments using different tuning methods across various models on testing split of Methods2Testsmall and HumanEval-Xjava datasets.
Model Method Trainable
params
Methods2Testsmall HumanEval-Xjava
Valid syntax
CodeBLEU
Valid syntax
CodeBLEU
pass@1
Instr Cov
Branch Cov
CodeGen-350M-multi None 0 95.43% 0.2170 100% 0.3608 0.0671 97.33% 89.77%
Full fine-tuning 304.23M 97.87% 0.2988 100% ( 0.3293 ) ( 0.0366 ) 100% ( 83.33% )
LoRA 1.31M ( 95.22% ) 0.2553 100% 0.3907 0.0671 98.69% ( 89.29% )
(IA)3 0.14M 95.53% 0.2266 100% ( 0.3583 ) ( 0.0549 ) 97.69% 94.32%
Prompt tuning 0.02M 96.03% 0.2208 100% ( 0.3290 ) ( 0.0427 ) ( 96.56% ) 91.67%
CodeGen2-1B None 0 0% 0 0% 0 0 0% 0%
Full fine-tuning 1,015.31M 76.73% 0.1474 5.49% 0.0359 0 0% 0%
LoRA 2.10M 41.16% 0.0484 8.54% 0.0117 0 0% 0%
(IA)3 0.23M 1.52% 0.2553 0% 0 0 0% 0%
Prompt tuning 0.04M 66.63% 0.2568 7.93% 0.2547 0 0% 0%
StarCoder2-3B None 0 85.23% 0.1543 100% 0.4264 0.3152 9.89% 85.67%
Full fine-tuning 3,030.37M 96.71% 0.2786 100% 0.4969 0.3494 99.40% 86.37%
LoRA 4.55M 97.11% 0.2901 100% ( 0.4169 ) 0.3675 99.19% 58.84%
(IA)3 0.47M 87.43% 0.2513 100% ( 0.4250 ) ( 0.2744 ) 99.67% ( 83.81% )
Prompt tuning 0.06M 86.43% 0.1742 100% 0.4309 ( 0.2470 ) 99.6% ( 75.85% )
CodeGen2-3.7B None 0 0% 0 0% 0 0 0% 0%
Full fine-tuning 3,641.17M 50.51% 0.1006 73.78% 0.2621 0 0% 0%
LoRA 4.19M 52.24% 0.0997 40.24% 0.1384 0 0% 0%
(IA)3 0.46M 0% 0 0% 0 0 0% 0%
Prompt tuning 0.08M 23.50% 0.2562 0% 0 0 0% 0%
CodeLlama-7B None 0 97.66% 0.3107 99.39% 0.4861 0.3293 98.33% 84.46%
Full fine-tuning 6,607.41M ( 96.44% ) ( 0.3012 ) 100% 0.4994 0.3373 98.95% 86.37%
LoRA 8.39M ( 97.36% ) 0.3277 99.39% ( 0.4291 ) ( 0.3129 ) 99.61% ( 72.47% )
(IA)3 0.61M ( 97.05% ) ( 0.3011 ) 100% ( 0.4802 ) ( 0.3232 ) 98.77% 84.72%
Prompt tuning 0.08M ( 95.93% ) ( 0.2885 ) 99.39% ( 0.4617 ) ( 0.2761 ) 98.38% ( 82.25% )
CodeGen2-7B None 0 96.95% 0.2848 100% 0.4736 0.2256 98.31% 81.45%
Full fine-tuning 6,862.87M 97.56% 0.3107 100% ( 0.4398 ) ( 0.1280 ) 99.75% ( 70.00% )
LoRA 8.39M 97.87% 0.3164 100% ( 0.4636 ) ( 0.2073 ) ( 98.06% ) ( 75.35% )
(IA)3 0.92M 97.36% 0.2904 100% 0.4898 ( 0.1829 ) 98.55% ( 79.50% )
Prompt tuning 0.08M ( 96.64% ) ( 0.2775 ) 100% ( 0.4407 ) ( 0.2012 ) 99.10% ( 69.40% )
StarCoder2-7B None 0 84.13% 0.1610 100% 0.4027 0.3758 99.07% 83.16%
Full fine-tuning 7,173.92M 97.21% 0.3009 100% 0.4389 ( 0.3675 ) 99.15% 90.80%
LoRA 7.34M 96.91% 0.3068 100% 0.5179 ( 0.3394 ) 99.35% 87.06%
(IA)3 0.75M 94.83% 0.2903 100% 0.4213 ( 0.3697 ) 99.40% 88.39%
Prompt tuning 0.09M ( 83.03% ) 0.3030 100% 0.5057 ( 0.3476 ) 99.38% 86.02%
StarCoderBase None 0 84.63% 0.1563 98.78% 0.4338 0.2963 99.07% 81.48%
Full fine-tuning 15,517.46M 96.81% 0.3123 100% 0.4830 0.3293 99.16% ( 75.20% )
LoRA 8.03M 95.71% 0.3152 98.78% ( 0.3905 ) 0.2963 99.07% ( 73.89% )
(IA)3 1.24M 84.63% ( 0.1553 ) 98.78% 0.4344 ( 0.1562 ) ( 98.72% ) 81.48%
Prompt tuning 0.12M 85.73% ( 0.1518 ) ( 78.05% ) ( 0.2315 ) 0.3025 99.76% ( 67.62% )
StarCoder2-15B None 0 85.43% 0.1898 100% 0.3724 0.4085 98.93% 87.69%
Full fine-tuning 15,655.90M 97.90% 0.3323 ( 99.39% ) 0.4886 0.3758 99.52% 81.3%
LoRA 12.12M 97.01% 0.3272 100% 0.4633 0.4146 98.95% 82.88%
(IA)3 1.25M 85.43% 0.1901 100% 0.3725 0.4578 99.57% 87.89%
Prompt tuning 0.12M 97.60% 0.3133 100% 0.5352 0.3939 99.32 % 82.89 %
CodeGen2-16B None 0 97.87% 0.2784 100% 0.4779 0.2012 98.66% 80.56%
Full fine-tuning 16,032.16M ( 97.56% ) 0.3383 ( 98.17% ) ( 0.3774 ) ( 0.1180 ) 99.52% ( 78.07% )
LoRA 13.37M 98.68% 0.3186 100% ( 0.4714 ) 0.2012 98.66% 82.06%
(IA)3 1.46M 97.87% 0.2790 100% 0.4780 0.2134 ( 97.46% ) 80.56%
Prompt tuning 0.08M 97.97% 0.2954 100% ( 0.4679 ) 0.2195 ( 98.62% ) ( 71.07% )
  • Bold: best-performing training method per model. ( Parentheses ): decreased performance compared to baseline. Red: $<$ 50% syntactical valid samples. Underline: Other notable results (see in Section V-A1).
V-A1 Syntactic Correctness
As shown in Table IV, the percentage of syntactically correctly generated unit tests is relatively high across the board, with $>$ 80% for all except the smaller CodeGen2-1B and CodeGen2-3.7B models. These two cannot produce any valid unit test with the base model. However, after tuning, both models are able to produce up to 76.73% and 52.24% syntactically valid tests, respectively. The notable findings (marked with underlines in Table IV) are as follows: 1) The increase of 66.63% using prompt tuning of CodeGen2-1B, tuning only 0.02‰ (0.04 million out of 1 billion parameters) of the parameters. 2) The increase of 52.24% using LoRA on CodeGen2-3.7B, tuning only 0.1% (4.19 million out of 3.7 billion parameters).
Further, data in Table IV illustrate that: 1) Full fine-tuning achieves higher percentages of syntactically correct unit tests than the baseline. The median percentage of syntactically valid tests for the full fine-tuning is 97.01%, which is 13.69% higher than the value of the baseline (85.33%). 2) There is a weak tendency that the more trainable parameters, the more samples are generated syntactically correctly. The median percentage of syntactically valid tests for the full fine-tuning is 97.01%, which is slightly higher than LoRA (96.96%), (IA)3 (91.13%), and prompt tuning (91.18%). However, considering that full fine-tuning often tunes parameters several thousands more times than PEFT methods, it is probably not wise to use full fine-tuning to improve the syntactic correctness of the unit tests if cost-effectiveness is a priority.
V-A2 Similarity
The Methods2Testsmall dataset CodeBLEU column in Table IV shows the CodeBLEU scores of the baseline, full fine-tuning, and PEFT methods. The notable results are: 1) LoRA archives the highest CodeBLEU scores in five of the ten models across different sizes. LoRA is also the only method that can positively increase the CodeBLEU score of the CodeLlama-7B model. 2) Full fine-tuning is nearly as performant as LoRA. For three of the ten models, full fine-tuning scores the best. 3) (IA)3 performs generally poorly compared to others, often resulting in very low CodeBLEU scores. 4) Prompt tuning shows varying results, from being the worst performing on CodeLlama-7B to being the best performing on CodeGen2-1B.
V-A3 Coverage
From the pass@1 column in Table IV, the average mean pass@1 rate is 0.2332. Of the tests that pass, the code coverage is generally high, as can be seen from the two last columns in Table IV. The exception is where the model is not able to generate much syntactically valid code. This results in some coverage results being zero, e.g. CodeGen2-1B and CodeGen2-3.7B. With an average median instruction coverage of 98.81% and branch coverage of 78.55%, this shows that if the models are able to create a runnable test, they are likely to cover a significant portion of the code. This high level of coverage indicates that the generated tests are effective at exercising the codebase, potentially identifying edge cases, and ensuring robustness. The results also show that LoRA and full fine-tuning outperform other methods in several cases. Although (IA)3 and prompt tuning perform best in some cases, the differences between their performance and others are insignificant.
V-A4 Resilience against catastrophic forgetting
The results in the last five columns of Table IV show the evaluation of various PEFT tuning methods on the Java subset of HumanEval-X after they have been trained on the Methods2Testsmall dataset. As expected, we can see from the results that there are some cases of degradation (marked using parentheses in Table IV) when comparing the results to the baseline. This applies to both full fine-tuning and PEFT. However, the levels of degradation are, for the most part, trivial. We also see several cases, e.g., StarCoder2-15B, where PEFT tuning actually improves the CodeBLEU scores. This shows that neither tuning method is very susceptible to catastrophic forgetting, and PEFT methods might pose similarly good transfer learning capabilities to that of full-fine-tuning. Summary for RQ1: Full fine-tuning and PEFT methods could improve the syntactic correctness of the generated unit tests. PEFT methods could be better choices than full fine-tuning if cost-effectiveness to improve the syntactic correctness of the generated unit tests is a concern. Full fine-tuning and LoRA could significantly boost the quality of the generated unit tests measured using CodeBLU, with LoRA even surpassing full fine-tuning in several cases. Further, PEFT methods are mostly resilient against catastrophic forgetting.
V-B RQ2: PEFT training resource utilization effectiveness
$10^{6}$$10^{8}$$10^{10}$rainable parametersCodeGen-350M-multiCodeGen2-1BStarCoder2-3BCodeGen2-3.7BCodeLlama-7BCodeGen2-7BStarCoder2-7BStarCoderBaseStarCoder2-15BCodeGen2-16BFull fine-tuningLoRA(IA)3Prompt tuning
Figure 2: Comparison of the number of trainable parameters for each tuning method. Models are sorted by size.
One of the main advantages of using PEFT methods is that they require significantly less computing during training, mainly because of reduced backpropagation calculations. The fewer trainable parameters, the more efficient. Results in Figure 2 show a significant reduction of trainable parameters of PEFT methods compared to full fine-tuning. This is also essential to fit big models into GPU memory. For example, we could train the biggest 16 billion parameter model using a single NVIDIA A100 80GB GPU for LoRA, whereas we needed 6 x NVIDIA A100 40GB GPU for full fine-tuning.
CodeGen-350M-multiCodeGen2-1BStarCoder2-3BCodeGen2-3.7BCodeLlama-7BCodeGen2-7BStarCoder2-7BStarCoderBaseStarCoder2-15BCodeGen2-16B1.01.0-1.0- 1.00.50.5-0.5- 0.50.00.00.00.00.50.50.50.51.01.01.01.0Relative importanceLoRA(IA)3Prompt tuning
Figure 3: Stacked bar chart of symmetrically normalized CodeBLEU score increase and decrease, relative to the fraction of trained parameters for each model. Positive values mean an increased CodeBLEU score, while negative values indicate a decreased CodeBLEU score.
Figure 3 shows the CodeBLEU score of each tuning method relative to the fraction of trainable parameters, symmetrically normalized between -1 and 1. As shown in Figure 3, prompt tuning has the highest efficiency among seven (i.e., CodeGen-350M-multi, CodeGen2-1B, StarCoder2-3B, CodeGen2-3.7B, StarCoder2-7B, StarCoder2-15B, and CodeGen2-16B) of the ten models and is the method that you get the ”most bang for your buck.” Moreover, prompt tuning has the most potential effect on models with large sizes, i.e., StarCoder2-15B and CodeGen2-16B. This can be explained by the increased capacity of larger models to understand better and leverage the additional context provided by prompt tuning. However, prompt tuning does not show consistent superiority for all the models. The method contributes to negative performance in three models (CodeLlama-7B, CodeGen2-7B, and StarCoderBase).
For the three smallest models (CodeGen-350M-multi, CodeGen2-1B, and StarCoder2-3B), (IA)3 clocks in at second place, which provides better results than LoRA. However, (IA)3 fails to improve the three largest models. LoRA achieves a relatively low efficiency compared to the other PEFT methods. However, it does not lead to a negative impact on the performance of all models. Full fine-tuning obviously has the worst efficiency, as it needs to train all the model parameters.
Summary for RQ2: PEFT methods offer significantly reduced computational requirements compared to full fine-tuning and are ranked by resource utilization effectiveness as follows: prompt tuning $>$ (IA)3 $>$ LoRA $>$ full fine-tuning. Prompt tuning demonstrates the highest returns on investment for most models but also carries risks of negative performance impact.
VI Discussion
This section highlights our novel insights by comparing our results with related work. We also discuss possible threats to the validity of our study and how we mitigate them.
VI-A Comparison with related works
Unlike existing studies, we investigated using PEFT and LLM to generate complete unit tests because the effectiveness of PEFT methods is task-dependent. Our results provide novel insights into their performance in the context of unit test generation. For example, Ayupov and Chirkova [5] conclude that LoRA is less effective than full fine-tuning for code generation. We show contradictory results that LoRA is more effective than fine-tuning in unit test generation. Weyssow et al. [8] explored generating code from natural language instructions and concluded that LoRA is the most effective one in the PEFT methods they evaluate. Our results show that prompt tuning is the most effective in terms of cost and resource utilization in unit test generation, although LoRA is more effective without considering the cost aspects.
Compared to the existing studies, we also performed a more comprehensive evaluation, as we invested significant efforts to compare all PEFT methods to full fine-tuning across all model sizes (up to 16B parameters). In contrast, Ding et al. [24] only evaluated PEFT using the 355 million parameter RoBERTaLARGE model, while Weyssow et al. [8] only did full fine-tuning for smaller models up to maximum 350 million parameters.
Several studies have leveraged LLMs for generating unit tests [1, 2, 3, 4, 29]. However, none of the works investigate the potential performance of using PEFT compared to full fine-tuning. Tufano et al. [1] automatically generated unit tests and reported 95% syntactic correctness when training BART large (400 million parameters) [52] on Methods2Test. With PEFT methods, we can generate an average median of 94.03% syntactically correct unit tests. Training between 0.02 million and 13 million parameters on 1% of the same data showcasing the potential powerful performance of PEFT methods. Our results also illustrate that LoRA surpasses full fine-tuning in terms of boosting the quality of the generated unit tests measured using CodeBLU in several cases.
VI-B Implications to Academia and Industry
Writing unit tests is a hard and often time-consuming task [53]. However, they are essential to ensure code quality, stability, and maintainability [54]. Automating unit test generation using LLMs can significantly reduce the time and effort required for this process, making it an attractive solution for academia and industry. However, as our experiments show, full fine-tuning is sometimes required to achieve adequate quality of the generated unit tests, which is computationally demanding. Many LLM users may not have the necessary computing resources to fine-tune LLMs fully. Our results show that various PEFT methods can achieve similar performance as full fine-tuning by training a minuscule parameter fraction. Furthermore, our results suggest starting with prompt tuning for cost-effectiveness. Then, if stability issues arise, LoRA should be considered as a reliable alternative because it does not lead to negative performance impact.
VI-C Threaths to Validity
VI-C1 External validity
One of the threats to external validity is the selection of models. We minimize this threat by selecting a broad range of models varying in size, pre-training data, and learning objectives. Another potential concern is the validity of the Methods2Testsmall benchmarking dataset. We opted for this dataset over other alternatives because it comprises real-world unit tests. We also validate the Java-translated version of the popularly used HumanEval code benchmarking dataset to mitigate the threat. The last threat is related to our evaluation metric. While popularly used, we recognize BLEU as a sub-optimal performance estimator for code generation. We, therefore, employ a more code-friendly version, CodeBLEU. Although CodeBLEU may not be the optimal metric to evaluate the functional correctness of unit tests completely, it provides a good tradeoff between evaluation effort and validity. To supplement our evaluation, we also calculate pass@1 and code coverage for the Methods2Testsmall dataset, as this contains isolated code samples that are easier to test.
VI-C2 Internal validity
The main threat to internal validity is comprised of the selection of hyperparameters. We use the hyperparameters reported in the introductory paper of each PEFT method. Additionally, we use greedy decoding to support reproducibility and reduce variability in the generated results.
VII Conclusion and future work
Although PEFT has been investigated to boost the quality and efficiency of code-related tasks using LLMs, they have not been studied in the context of unit test generation. We performed the first empirical study to compare different PEFT methods and full fine-tuning to generate unit tests. Our comparison involved ten LLMs of different sizes and architectures. Our results offer valuable insights into the advantages and challenges of using PEFT and full fine-tuning for unit test generation. With the knowledge we’ve gathered, individuals can choose the most suitable PEFT methods or full fine-tuning based on the size and characteristics of the LLMs.
In addition to unit tests, manually coding other automated test cases, such as integration and system tests, can be cost-intensive. In the future, we intend to explore the utilization of LLMs in conjunction with cost-effective tuning methods to automatically generate other types of tests.
VIII Data availability
The Methods2Testsmall dataset is available at:
https://huggingface.co/datasets/andstor/methods2test_small
All fine-tuned models are also made available at:
https://huggingface.co/andstor/peft-unit-test-generation-experiments
Experiment code and results are available at:
https://github.com/andstor/peft-unit-test-generation-replication-package

### v2

@@
 Abstract
-The advent of large language models (LLMs) like GitHub Copilot has significantly enhanced programmers’ productivity, particularly in code generation.
-However, these models often struggle with real-world tasks without fine-tuning.
-As LLMs grow larger and more performant, fine-tuning for specialized tasks becomes increasingly expensive.
-Parameter-efficient fine-tuning (PEFT) methods, which fine-tune only a subset of model parameters, offer a promising solution by reducing the computational costs of tuning LLMs while maintaining their performance.
+Parameter-efficient fine-tuning (PEFT) methods, which fine-tune only a subset of model parameters, offer a promising solution by reducing the computational costs of tuning large language models (LLMs) while maintaining their performance.
 Existing studies have explored using PEFT and LLMs for various code-related tasks and found that the effectiveness of PEFT techniques is task-dependent.
+The state-of-the-art is limited to using LLMs with full fine-tuning to generate unit tests.
 The application of PEFT techniques in unit test generation remains underexplored.
-The state-of-the-art is limited to using LLMs with full fine-tuning to generate unit tests.
-This paper investigates both full fine-tuning and various PEFT methods, including LoRA, (IA)3, and prompt tuning, across different model architectures and sizes.
-We use well-established benchmark datasets to evaluate their effectiveness in unit test generation.
-Our findings show that PEFT methods can deliver performance comparable to full fine-tuning for unit test generation, making specialized fine-tuning more accessible and cost-effective.
-Notably, prompt tuning is the most effective in terms of cost and resource utilization, while LoRA approaches the effectiveness of full fine-tuning in several cases.
-Index Terms:
-large language models, parameter-efficient fine-tuning, test generation, unit tests
-I Introduction
-Large language models (LLMs) have been developed for various tasks, including general language and code-related activities, significantly enhancing programmers’ performance.
-Training LLMs on extensive code bases like GitHub give them seemingly magical capabilities.
-However, despite their proficiency in multiple code-related tasks, LLMs still struggle with cost-effectiveness in their training.
-The traditional approach to handling these specialized tasks has been fine-tuning the model specifically for each task.
-Given that LLMs often have billions, if not trillions, of parameters, traditional fine-tuning, which adjusts all parameters, is extremely computationally expensive and resource-intensive.
-However, state-of-the-art approaches [1] still rely on energy-costly full fine-tuning.
-Advancements in instruction-tuned LLMs have encouraged investigations into various prompting and tuning techniques, e.g., [2, 3, 4], that do not require fine-tuning.
+This paper investigates both full fine-tuning and various PEFT methods, including LoRA, (IA)3, and prompt tuning, across thirteen models of different architectures and sizes.
+We use well-established benchmark datasets to evaluate their effectiveness in unit test generation and measure syntax correctness, CodeBLEU, pass@1, instruction coverage, branch coverage, and mutation score of the generated tests.
+Our findings show that LoRA can deliver performance comparable to full fine-tuning for unit test generation in several cases.
+If training costs are valued, prompt tuning is the most cost-effective approach, particularly for large models.
+However, the models tuned with full fine-tuning or PEFT may generate fewer executable test cases than the baseline model because they generate more tests calling nonexistent methods or having type mismatches.
+For the generated ones that are executable, the ones from the tuned models show better test coverage than those from the baseline model.
+1 Introduction
+Training LLMs on extensive code bases like GitHub gives them seemingly magical capabilities.
+However, despite their proficiency in multiple code-related tasks, LLMs still struggle with the high cost of their training.
+The traditional approach to handling specialized tasks has been fine-tuning the model for each task.
+Given that LLMs often have billions, if not trillions, of parameters, traditional fine-tuning, which adjusts all parameters, is extremely computationally expensive and resource-intensive [tufano2021unit].
+Advancements in LLMs have encouraged investigations into various prompting and tuning techniques, e.g., [siddiq2024using, yuan2023manual, schäfer2023empirical], that do not require full fine-tuning.
 This challenge has led to the development of multiple techniques for fine-tuning larger models with limited resources, commonly referred to as parameter-efficient fine-tuning (PEFT).
-The application of various PEFT methods has been explored for code-related generation tasks, such as code completion, code summarization, and more [5, 6, 7, 8, 9, 10, 11, 12, 13].
-Evaluations of the PEFT methods show that they could achieve performance comparable to full fine-tuning while significantly reducing the computational burden by strategically fine-tuning only a chosen subset of parameters during the code-related generation process.
-However, the effectiveness of PEFT techniques on different tasks varies.
-LLMs have been used to generate unit test cases and assertion statements [14, 12, 2, 3, 4, 1].
-However, to the best of our knowledge, the area of unit test generation using PEFT remains largely unexplored.
+The application of various PEFT methods has been explored for code-related tasks, such as code completion, code summarization, and more [Ayupov2022ParameterEfficientFO, Goel2022OnTC, Wang2023PromptTI, weyssow2024exploring, Wang2023OneAF, Shi2023TowardsEF, Liu2023MFTCoderBC, yuan2023evaluating, Liu2023AnES].
+Evaluations of the PEFT methods show that they could achieve performance comparable to full fine-tuning while significantly reducing the computational burden by strategically fine-tuning only a chosen subset of parameters for code-related tasks.
+However, the effectiveness of PEFT techniques on different tasks varies [Ding2023ParameterefficientFO].
+LLMs have been used to generate unit test cases and assertion statements [Mastropaolo2021StudyingTU, yuan2023evaluating, siddiq2024using, yuan2023manual, schäfer2023empirical, tufano2021unit].
+However, to the best of our knowledge, the area of empirical comparisons of PEFT and full fine-tuning for unit test generation remain unexplored.
 Therefore, in this work, we empirically evaluate the performance of various PEFT methods for unit test generation.
@@
   • RQ2: What is the relation between resource utilization and performance of PEFT in unit test generation?
-To answer the research questions, we compared full fine-tuning with three popular PEFT methods: LoRA (Low-Rank Adaptation) [15], (IA)3 (Infused Adapter by Inhibiting and Amplifying Inner Activations) [16], and prompt tuning [17].
-We conducted these comparisons using ten LLMs from three open-source, decoder-only LLM families.
+To answer the research questions, we compared full fine-tuning with three popular PEFT methods: LoRA (Low-Rank Adaptation) [Yu2023LowRankAO], (IA)3 (Infused Adapter by Inhibiting and Amplifying Inner Activations) [Liu2022FewShotPF], and prompt tuning [Lester2021ThePO].
+We conducted the comparisons using thirteen LLMs from four open-source decoder-only LLM families.
 The LLMs have varying architectures and sizes, ranging from 350 million to 16 billion parameters.
-We measured and compared the unit test generation performance of full fine-tuning and PEFT methods using well-established benchmark datasets rooted in real-world unit test cases.
-Our results show that LoRA is generally the most reliable method for improving the quality of the generated unit tests and is, in many cases, on par with or better than full fine-tuning.
-Conversely, (IA)3 appears to be the least effective, often resulting in minimal improvements while consistently retaining prior knowledge in the LLMs.
-Prompt tuning demonstrates significant variability, excelling with some models, particularly larger ones, while underperforming with others.
-This variability highlights the importance of considering model size and characteristics when choosing the appropriate tuning method.
-Prompt tuning can be considered ”low-hanging fruit,” while LoRA should be employed for greater stability.
+We measured and compared the unit test generation performance of full fine-tuning and PEFT methods using well-established benchmark datasets rooted in real-world unit test cases and metrics, including syntactic correctness, CodeBLEU, pass@1, instruction coverage, branch coverage, and mutation score.
+Our results show that LoRA is generally the most reliable PEFT method to generate good-quality unit tests and is, in many cases, on par with or even better than full fine-tuning.
+Regarding cost-effectiveness, prompt tuning provides the best results, particularly for larger models.
+Prompt tuning can be considered “low-hanging fruit”, while LoRA should be employed for greater performance.
+The results also reveal that the generated unit tests from tuned models can contain more compilation errors, which need to be investigated in future studies.
 Our contributions are as follows:
-  • We conducted the first empirical study to extensively evaluate training LLMs using various PEFT methods—LoRA, (IA)3, and prompt tuning—for unit test generation across a wide range of LLM models.
+  • We conducted the first empirical study to extensively evaluate tuning LLMs using various PEFT methods—LoRA, (IA)3, and prompt tuning—for unit test generation across a wide range of LLM models.
   • We offered a comprehensive comparison of PEFT methods versus full fine-tuning.
@@
 The rest of the paper is organized as follows.
-We provide preliminaries in Section II.
-Section III introduces related work.
-We explain our experimental design in Section IV.
-Section V presents the experimental results.
-Section VI discusses our results and limitations.
-Section VII concludes the study and proposes future work.
-II Preliminaries
-II-A LLM training
-The classic language modeling task involves predicting the next word in a sequence, given the preceding words.
-The model is trained in an autoregressive manner.
-A typical example is the GPT model [18].
-Training an autoregressive model involves fitting a function $f:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$, where $N$ represents the sequence length of tokens, each token having a dimension of $D$.
-The model aims to estimate the conditional probability distribution, denoted $p$, which can be expressed as:
-$p(x_{1:T}\mid c)=\prod_{t=1}^{T}p(x_{t}\mid x_{1:t-1},c)$ (1)
-where $x_{t}\in\mathbb{R}^{D}$ represents the $t$’th discrete token observation, $c$ denotes the initial context or input sequence, and $T$ signifies the length of the sequence of predicted tokens [19, Chapter 22].
-LLMs are normally pre-trained on a huge amount of data, giving them a broad understanding of language patterns and semantics.
-This pre-training phase allows LLMs to capture intricate linguistic structures and contextual nuances, empowering them with a robust foundation for various natural language processing tasks.
-Fine-tuning is a method in which a pre-trained model can be adapted to a more specific task, such as auto-completing source code.
-The process involves adjusting the weights of the pre-trained model using task-specific data, thereby enhancing its performance on a target task.
-When all parameters are tuned, this is often referred to as “full fine-tuning” [20].
-If we imagine the model as a web of knowledge, the full fine-tuning loosens all the connections and lets them adapt entirely to the new domain.
-This flexibility allows for potentially superior performance but comes at a cost:
-  • Overfitting: The model can become overly reliant on the training data, potentially struggling with unseen examples [21].
-  • Computational cost: Training all parameters requires significant computational resources [20].
-  • Catastrophic forgetting: The model might forget some valuable knowledge it acquired during pre-training [22, 23].
-II-B PEFT Methods
-Parameter-efficient fine-tuning (PEFT) methods provide an alternative way of fine-tuning a pre-trained model by training so-called adapters [24].
+Section 2 introduces the PEFT methods we explore.
+Section 3 introduces related work.
+We explain our experimental design in Section 4.
+Section 5 presents the experimental results.
+Section 6 discusses our results and limitations.
+Section 7 concludes the study and proposes future work.
+2 PEFT Methods
+Parameter-efficient fine-tuning (PEFT) methods provide an alternative way of fine-tuning a pre-trained model by training so-called adapters [Ding2023ParameterefficientFO].
 Here, most (if not all) parameters of the pre-trained model are frozen, and only a small number of parameters are trained.
 These adapters are often orders of magnitude smaller than the full model, making sharing, storing, and loading especially convenient.
-Moreover, some PEFT methods even facilitate the utilization of mixed-task batches, enabling distinct processing for different examples within a batch [17].
-PEFT methods can be classified into three main classes of methods: additive, selective, and reparametrization-based [25].
-Additive methods can be further broken down into adapters  [^1] The term ”adapter” is also used as the name of the first proposed adapter module for NLP.
+Moreover, some PEFT methods even facilitate the utilization of mixed-task batches, enabling distinct processing for different examples within a batch [Lester2021ThePO].
+PEFT methods can be classified into three main categories: additive, selective, and reparametrization-based [lialin2023scaling].
+Additive methods can be further broken down into adapters  [^1] The term “adapter” is also used as the name of the first proposed adapter module for NLP.
 and soft prompts.
-II-B1 Additive
+2.1 Additive
 This approach focuses on augmenting the pre-trained model with additional parameters or layers.
 Only the newly added components are trained, keeping the original model weights frozen.
-Popular techniques within the additive category are (IA)3 [16], adapters [26], and prompt tuning [17].
+Popular techniques within the additive category are (IA)3 [Liu2022FewShotPF], adapters [houlsby2019parameter], and prompt tuning [Lester2021ThePO].
 Adapters
-Adapter, or “bottleneck adapter,” introduced by Houlsby et al.
-[26] was arguably the first PEFT method.
+Adapter, or “bottleneck adapter,” introduced by [houlsby2019parameter], was arguably the first PEFT method.
 The class of “adapters” introduces small, fully connected networks placed after specific sub-layers within the model.
@@
 Soft prompts
-These techniques focus on creating the so-called “soft prompts.” Prompt tuning, introduced by Lester et al.
-[17], fine-tuned a subset of the model’s input embeddings, allowing the model to adapt to new tasks without altering the model’s parameters.
-The prefix tuning expands the idea to all model layers [27].
-II-B2 Selective
+These techniques focus on creating the so-called “soft prompts.” Prompt tuning, introduced by [Lester2021ThePO], fine-tuned a subset of the model’s input embeddings, allowing the model to adapt to new tasks without altering the model’s parameters.
+The prefix tuning expands the idea to all model layers [li-liang-2021-prefix].
+2.2 Selective
 When using selective approaches, only a specific subset of the pre-trained model’s parameters is fine-tuned.
 This approach requires carefully selecting the parameters to be updated, ensuring that they are most relevant to the new task.
-An example is BitFit by Ben Zaken et al.
-[28], which only updates the bias parameters.
-II-B3 Reparametrization-based
+An example is BitFit by [BenZaken2021BitFitSP], which only updates the bias parameters.
+2.3 Reparametrization-based
 This category utilizes techniques that reformulate the original model’s parameters into a more efficient trainable representation.
 This makes it possible to work with high-dimensional matrices while simultaneously reducing the number of trainable parameters.
-LoRA [15] is a common reparametrization-based method.
-III Related Work
-This section presents state-of-the-art studies generating unit tests using LLMs and PEFT approaches to enhance code-related tasks.
-III-A Unit test generation using LLMs
-Researchers have turned to neural techniques to automatically generate unit tests to align with real-world best practices in unit test generation.
-Tufano et al.
-[1] propose an automated unit test case generation tool called AthenaTest.
-They construct a real-world focal method and developer-written test case dataset named Methods2Test 222 https://github.com/microsoft/methods2test and train a sequence-to-sequence transformer to generate unit test cases.
-Advancements in instruction-tuned LLMs have also encouraged several investigations of various prompting techniques, as these do not require any fine-tuning [2, 3, 4].
-Yuan et al.
-[3] propose ChatTester, a ChatGPT-based unit test generation approach, which leverages ChatGPT itself to improve the quality of its generated tests.
-Schäfer et al.
-[4] conducted a large-scale empirical evaluation on the effectiveness of LLMs for automated unit test generation without additional training or manual effort.
-Similarly to ChatTester, they also use ChatGPT (GPT-3.5-Turbo) and propose a re-prompting approach named TestPilot that automatically generates unit tests for all API functions in an npm package.
-Siddiq et al.
-[2] use the HumanEval and SF110 datasets to benchmark how well Codex, GPT-3.5-Turbo, and StarCoder can automatically generate unit tests.
+LoRA [Yu2023LowRankAO] is a typical reparametrization-based method.
+3 Related Work
+This section presents state-of-the-art studies generating unit tests using LLMs and studies applying PEFT approaches to enhance code-related tasks.
+3.1 Unit test generation using LLMs
+[tufano2021unit] propose an automated unit test case generation tool called AthenaTest.
+They construct a real-world focal method and developer-written test case dataset named Methods2Test  [^2] https://github.com/microsoft/methods2test and train a sequence-to-sequence transformer to generate unit test cases.
+[yuan2023manual] propose ChatTester, a ChatGPT-based unit test generation approach, which leverages ChatGPT itself to improve the quality of its generated tests.
+[schäfer2023empirical] conducted a large-scale empirical evaluation on the effectiveness of LLMs for automated unit test generation without additional training.
+Similarly to ChatTester, they also use ChatGPT (GPT-3.5-Turbo) and propose a re-prompting approach named TestPilot.
+TestPilot automatically generates unit tests for all API functions in an npm package, achieving a median statement and branch coverage of 70.2% and 52.8% on 1,684 API functions, respectively.
+[siddiq2024using] use the HumanEval and SF110 datasets to benchmark how well Codex, GPT-3.5-Turbo, and StarCoder can automatically generate unit tests.
 They evaluate the models based on compilation rates, test correctness, test coverage, and test smells.
-III-B PEFT methods on code-related tasks
-While prompting methods are gaining traction, they often require significant manual effort to design complex templates and rely on expensive enterprise-grade LLMs like ChatGPT.
-Consequently, most cutting-edge methods for automatic code generation still involve computationally expensive full fine-tuning.
-One alternative approach to full fine-tuning is proposed by Steenhoek et al.
-[29].
+[alagarsamy2024a3test] propose A3Test, a deep-learning based unit-test generation method that augments test generation with assertion knowledge and verifies naming and signatures.
+They evaluated the method on 5,278 focal methods from the Defects4J [just2014defects4j] benchmark.
+The methods by [yuan2023manual, siddiq2024using, schäfer2023empirical, alagarsamy2024a3test] primarily leverage large off-the-shelf LLMs, which can be costly and often lack the flexibility for custom fine-tuning to generate unit tests for specific domains and organizations.
+Additionally, such prompting-based methods often require significant manual effort to design complex templates and rely on expensive enterprise-grade LLMs like ChatGPT.
+3.2 PEFT methods on code-related tasks
+Many cutting-edge methods for automatic code generation still involve computationally expensive full fine-tuning.
+One alternative approach to full fine-tuning is proposed by [steenhoek2023reinforcement].
 They use Reinforcement Learning from Static Quality Metrics (RLSQM).
@@
 PEFT has emerged as a powerful technique for enhancing code generation with LLMs.
-Researchers are investigating how these models perform on code tasks using PEFT techniques as a more efficient alternative to full fine-tuning and PEFT techniques’ effectiveness for different code-related tasks.
-Results show that the effectiveness of PEFT techniques is task-dependent and not transferrable.
-For example, Ayupov and Chirkova [5] evaluated adapters and LoRA on various tasks, finding them effective for code understanding but less so for code generation compared to full fine-tuning.
-Wang et al.
-[7] reports that prompt tuning of CodeBERT and CodeT5 outperforms full fine-tuning for defect prediction, code summarization, and code translation tasks.
-Liu et al.
-[13] evaluated adapter, LoRA, Prefix tuning, and Mix-And-Match adapter (MAM) [30] for defect detection, code clone detection, and code summarization.
-They concluded that “PEFT may achieve comparable or higher performance than full fine-tuning in code understanding tasks, but they may exhibit slightly weaker performance in code generation tasks.” Similarly, Weyssow et al.
-[8] evaluate text-to-code generation using LoRA, (IA)3, prompt tuning, and prefix tuning.
-They compare performance across models of varying sizes (up to 7 billion parameters) against full fine-tuning of smaller models (less than 350 million parameters) and concluded that LoRA is the most effective one.
-To our knowledge, there is currently no systematic knowledge of how various PEFT methods behave in the domain of unit test generation.
+Researchers are investigating how to apply PEFT techniques to code-related tasks as a more efficient alternative to full fine-tuning.
+Results show that the effectiveness of PEFT techniques is task-dependent and not transferable.
+For example, [Ayupov2022ParameterEfficientFO] evaluated adapters and LoRA on various tasks, finding them effective for code understanding but less so for code generation compared to full fine-tuning.
+[Wang2023PromptTI] report that prompt tuning of CodeBERT and CodeT5 outperforms full fine-tuning for defect prediction, code summarization, and code translation tasks.
+[Liu2023AnES] evaluated adapter, LoRA, Prefix tuning, and Mix-And-Match adapter (MAM) [mam] for defect detection, code clone detection, and code summarization.
+They concluded that “PEFT may achieve comparable or higher performance than full fine-tuning in code understanding tasks, but they may exhibit slightly weaker performance in code generation tasks.” Similarly, [weyssow2024exploring] evaluated text-to-code generation using LoRA, (IA)3, prompt tuning, and prefix tuning.
+They compared performance across models of varying sizes (up to 7 billion parameters) against full fine-tuning of smaller models (less than 350 million parameters) and concluded that LoRA is the most effective one.
 As existing studies show that PEFT techniques’ effectiveness is task-dependent, it is, therefore, necessary to empirically evaluate their effectiveness for unit test generation.
-IV Experimental design
-The main objective of this study is to gain insight into the performance of fine-tuning LLMs using PEFT for unit test generation.
-IV-A Research Questions
-To address RQ1 (”How well does PEFT perform on unit test generation?”), we evaluate several different PEFT methods on unit test generation.
-We also compare the PEFT methods with full fine-tuning.
-Our study covers a wide range of language models for code of various sizes, spanning 350 million to 16 billion parameters.
-We assess the quality of the generated test cases produced by different PEFT methods by scoring against real-world human-written reference tests and evaluating the test cases’ syntactic correctness.
-Further, we examine whether applying PEFT methods for unit test generation adversely affects the pre-trained models’ code generation performance, also known as catastrophic forgetting [22, 23].
+To our knowledge, there is currently no systematic knowledge of how various PEFT methods behave to tune LLMs for unit test generation.
+In this study, we focus on autoregressive base code models to minimize the influence of prompt templates on PEFT performance.
+4 Experimental design
+The main objective of this study is to gain insight into the performance and cost-effectiveness of fine-tuning LLMs using PEFT for unit test generation.
+4.1 Research Questions
+To answer RQ1 (“How well does PEFT perform on unit test generation?”), we aim to evaluate several different PEFT methods on unit test generation.
+We also want to compare the PEFT methods with full fine-tuning.
+Thus, our study covers a wide range of language models for code of various sizes, spanning 350 million to 16 billion parameters.
+We assess the quality of the generated test cases by different PEFT methods by scoring against real-world human-written reference tests and evaluating the test cases’ syntactic correctness, similarity to reference tests, passing rates, instruction and branch coverages, and mutation scores.
+Further, we examine whether applying PEFT methods for unit test generation adversely affects the pre-trained models’ code generation performance, also known as catastrophic forgetting [French1999CatastrophicFI, Luo2023AnES].
+Organizations may use the same model to generate both code and unit tests within their integrated development environments (IDEs).
 This evaluation will help us understand if there are any trade-offs between applying fine-tuning and PEFT methods for unit test generation and maintaining performance on other code-related tasks.
-Regarding RQ2 (”What is the relation between resource utilization and performance of PEFT in unit test generation?”), we want to understand PEFT methods’ performance gains in relation to their resource utilization and requirements.
-Thus, we analyze the efficiency of PEFT methods in terms of the computational resources they use and their impact on the performance outcomes in unit test generation.
-IV-B Fine-tuning methods selection
-In line with other empirical studies on PEFT [8, 31], we select the popularly used PEFT methods: LoRA [15], (IA)3 [16], and prompt tuning [17].
+To answer RQ2 (“What is the relation between resource utilization and performance of PEFT in unit test generation?”), we analyze the efficiency of PEFT methods in terms of the computational resources they use and the performance they achieve in unit test generation.
+4.2 Fine-tuning methods selection
+In line with other empirical studies on PEFT [weyssow2024exploring, Pu2023EmpiricalAO], we select the most popularly used PEFT methods: LoRA [Yu2023LowRankAO], (IA)3 [Liu2022FewShotPF], and prompt tuning [Lester2021ThePO].
 They are readily available in popular libraries and exemplify most of the primary approaches within PEFT.
-(a) Diagram of LoRA (based on [15]).
-(b) Diagram of (IA)3 (based on [16]).
-(c) Diagram of prompt tuning (based on [17]).
+(a) Diagram of LoRA (based on [Yu2023LowRankAO]).
+(b) Diagram of (IA)3 (based on [Liu2022FewShotPF]).
+(c) Diagram of prompt tuning (based on [Lester2021ThePO]).
 Figure 1: PEFT methods architecture illustration diagrams based on the reference implementation of each introductory paper. Striped modules are frozen. Magenta modules are trained.
-  • LoRA: A reparametrization-based method that updates a weight matrix by training a new matrix that is decomposed into a product of two matrices of lower rank: $\delta W=W_{A}\cdot W_{B}$, where $W_{A}\in\mathbb{R}^{\text{in}\times r}$ and $W_{B}\in\mathbb{R}^{r\times\text{out}}$ [15].
+  • LoRA: A reparametrization-based method that updates a weight matrix by training a new matrix that is decomposed into a product of two matrices of lower rank: $\delta W=W_{A}\cdot W_{B}$, where $W_{A}\in\mathbb{R}^{\text{in}\times r}$ and $W_{B}\in\mathbb{R}^{r\times\text{out}}$ [Yu2023LowRankAO].
 As demonstrated in Figure 1(a), the decomposed matrix is then added to the original frozen K and V projection matrices in the attention modules.
 All other modules are kept frozen.
-  • (IA)3: An additive method that focuses on fine-tuning specific intermediate activations within the model [16].
+  • (IA)3: An additive method that focuses on fine-tuning specific intermediate activations within the model [Liu2022FewShotPF].
 As illustrated in Figure 1(b), this is done by learning three new vectors: $l_{v},l_{k},l_{ff}$; rescaling key, value, and hidden feed-forward activations, respectively.
-  • Prompt tuning: A soft prompts method that fine-tunes a subset of the model’s input embeddings.
-Exemplified in Figure 1(c), the model input embeddings are prepended with a trainable tensor $P\in\mathbb{R}^{N\times D}$ (known as a “soft prompt”) [17].
+  • Prompt tuning: A soft prompts (additive) method that fine-tunes a subset of the model’s input embeddings.
+Exemplified in Figure 1(c), the model input embeddings are prepended with a trainable tensor $P\in\mathbb{R}^{N\times D}$ (known as a “soft prompt”) [Lester2021ThePO].
 Each element of $P$ is also referred to as a “virtual token.” During training, these “virtual tokens” are updated while all the rest of the model parameters are kept frozen.
-IV-C Models
-For our experiments, we select a diverse set of open-source decoder-only model families trained on code to evaluate the performance and efficiency of PEFT methods.
-The chosen models span a range of sizes and architectures and are trained on diverse datasets and training objectives.
+4.3 Models
+We select a diverse set of open-source autoregressive model families trained on code to evaluate the performance and cost-efficiency of PEFT methods.
+The chosen models span a range of sizes and architectures from the last four years, trained on diverse datasets and training objectives.
 This offers extensive insights into the interplay between various model capacities and PEFT methods.
-Further, all models are ensured to support the Java programming language to facilitate a fair comparison of our benchmarking datasets described in Section IV-D.
-  • Salesforce CodeGen Models: To democratize program synthesis, Salesforce AI Research developed CodeGen [32].
+Further, all models are ensured to support the Java programming language to facilitate a fair comparison of our benchmarking datasets described in Section 4.4.
+  • Salesforce CodeGen Models: To democratize program synthesis, Salesforce AI Research developed CodeGen [nijkamp2022codegen].
 Pioneering in multi-turn program synthesis, CodeGen presents a variety of sizes, ranging from 350 million to 16 billion parameters.
 These models have undergone training across a diverse array of both programming and natural languages.
-The models were trained on The pile [33], followed by a code subset of Google’s BigQuery datasets Cloud [34].
-An updated version CodeGen2 [35] is also available in sizes 1 billion to 16 billion, trained on The Stack dataset [36] using both causal language modeling (CLM) and file-level span corruption.
+The models were trained on “The pile” [gao2020pile], followed by a code subset of Google’s BigQuery datasets [GoogleBigQuery].
+An updated version CodeGen2 [nijkamp2023codegen2] is also available in sizes 1 billion to 16 billion, trained on “The Stack” dataset [Kocetkov2022TheStack] using both causal language modeling (CLM) and file-level span corruption.
 We select the smallest model of the first generation CodeGen family (CodeGen-350M-multi), along with all model sizes of the second generation (1 billion, 3.7 billion, 7 billion, and 16 billion parameters).
-  • Meta LLaMA Models: Code Llama [37], derived from Llama 2 [38], is a specialized model further fine-tuned on the code-specific parts of its pre-training data.
+  • Meta LLaMA Models: Code Llama [Rozire2023CodeLO], derived from Llama 2 [Touvron2023Llama2O], is a specialized model further fine-tuned on the code-specific parts of its pre-training data.
 With enhanced coding capabilities, it can generate code, interpret natural language prompts, assist in code completion, and aid debugging.
@@
 We select the base variant CodeLlama-7B, as it is designed for general code synthesis and understanding and supports a wide range of popular languages, including Java.
-  • BigCode StarCoder Models: This model family includes the 15.5 billion parameters StarCoderBase model [39], along with an updated model generation named StarCoder2 [40], available in 3 billion, 7 billion, and 15 billion parameters.
-StarCoderBase was trained on 80+ programming languages from The Stack dataset [36].
+  • BigCode StarCoder Models: This model family includes the 15.5 billion parameters StarCoderBase model [li2023starcoder], along with an updated model generation named StarCoder2 [Lozhkov2024StarCoder2A], available in 3 billion, 7 billion, and 15 billion parameters.
+StarCoderBase was trained on 80+ programming languages from “The Stack” dataset [Kocetkov2022TheStack].
 The updated StarCoder2 is trained on 619 programming languages and other data sources, such as GitHub pull requests, Kaggle notebooks, and code docs.
 We select all model generations and sizes in this model family.
-IV-D Datasets
-IV-D1 Methods2Test
-To ensure our experiments’ validity and practical applicability, we rely on Methods2Test [1].
+  • Qwen 2.5 Coder Models: Qwen 2.5 Coder [hui2024qwen] is the latest code-specific series of Qwen LLMs.
+The series comes in 0.5, 1.5, 3, 7, 14, 32 billion parameters.
+The models are built upon the Qwen2.5 architecture and pre- trained on a vast corpus of over 5.5 trillion tokens.
+The dataset is sourced from a wide range of public code repositories, as well as large-scale web-crawled data containing code-related texts.
+We select the 3, 7 and 14 bllion parameter sizes of this model family.
+Table 1 shows model statistics for the selected models derived from cfahlgren1/hub-stats hugging face dataset [^3] https://huggingface.co/datasets/cfahlgren1/hub-stats as of 13 November 2025.
+The chosen models span a representative set of model families released over the past four years, many of which remain highly popular.
+Table 1: HuggingFace model download statistics from cfahlgren1/hub-stats dataset as of 13 November 2025.
+Model Downloads in October 2025 Downloads all time Created at
+Qwen/Qwen2.5-Coder-14B 4621 185216 2024-11-08
+Qwen/Qwen2.5-Coder-3B 15469 213837 2024-11-08
+Qwen/Qwen2.5-Coder-7B 22839 453481 2024-09-16
+meta-llama/CodeLlama-7b-hf 2274 200519 2024-03-13
+bigcode/starcoder2-7b 17013 695181 2024-02-20
+bigcode/starcoder2-15b 5397 450456 2024-02-20
+bigcode/starcoder2-3b 192177 6168133 2023-11-29
+bigcode/starcoderbase 7185 210970 2023-05-03
+Salesforce/codegen2-16B_P 13 37908 2023-04-26
+Salesforce/codegen2-7B_P 90 28639 2023-04-26
+Salesforce/codegen2-3_7B_P 42 39302 2023-04-25
+Salesforce/codegen2-1B_P 226 77365 2023-04-25
+Salesforce/codegen-350M-multi 4113 588888 2022-04-11
+4.4 Datasets
+4.4.1 Methods2Testsmall
+To ensure our experiments’ validity and practical applicability, we rely on Methods2Test by [tufano2021unit].
 It is the largest publicly available corpus of Java unit test cases and corresponding focal methods.
@@
 Thus, we opt to downsize the Methods2Test dataset to enhance efficiency without compromising the representativeness of the data.
-Specifically, we adopt a strategy wherein only one focal function is selected from each class in the dataset.
-By doing so, we maintain the essence and diversity of the original dataset while significantly reducing its size.
+Specifically, we adopt a strategy wherein only one focal function is selected at random from each class in the dataset.
+By doing so, we maintain the essence and diversity of the original dataset while significantly reduce its size.
 Further, the Methods2Test dataset is distributed in deconstructed components.
-Hence, we assemble it according to the structure seen in LABEL:lst:data-structure and repackage it as Methods2Testsmall [^3] Available at https://huggingface.co/datasets/andstor/methods2test_small.
-The resulting dataset contains 7,44K training, 953 validation, and 1.02K testing tests.
+Hence, we assemble it according to the structure seen in LABEL:lst:data-structure and package it as Methods2Testsmall.
+The resulting dataset contains 7440 training, 953 validation, and 1017 testing tests.
+We use this as the primary training dataset.
+4.4.2 Methods2Testrunnable
+As the Methods2Test dataset is made up of tests from the real world, executing the tests to measure their passing rates, test coverages, and mutation scores needs significant extra work.
+This includes retrieving the correct repository version, resolving dependencies, injecting coverage collection tools and mutation testing systems into the existing build system, and finally executing the code in a controlled and safe manner.
+We construct a runnable subset of Methods2Test by the following steps: 1) Golden commit: Methods2Test does not provide the repository commits at the time of scraping.
+Therefore, for each repository, we consider all commits prior to the last commit [^4] https://github.com/microsoft/methods2test/commit/309ee31cb3752cd8f3ad4b33cfd2b6e5b0ec5029 on GitHub that updated the Methods2Test dataset (2021-05-18T23:18:51.000Z UTC) in which the focal method and test class files exist as candidate commits.
+We brute force (begin at the most recent candidate commit and iteratively check older commits) compare all available dataset data against the repository until we find an exact match—a golden commit.
+2) Build system: We target repositories using the Maven build system [apache-maven:3.8.1], as this is somewhat structured and not as flexible as Gradle.
+We rely on build file names to determine the build system.
+Of the 1017 tests in the test split of Methods2Testsmall, 624 of the repositories in Methods2Test are using Maven, 264 are using Gradle, and 129 are using others.
+3) Runnable: Each repository in Methods2Test’s test split is then checked for buildability by running Maven and executing the test split of Methods2Testsmall.
+Due to a lot of old and missing dependencies, only 119 of the 624 Maven repositories are buildable.
+We collect all tests from Methods2Test with buildable repositories, resulting in 6150 tests.
+We then validate their runnability by executing every test, obtaining 3712 runnable tests (see section 4.8 for additional test execution details.) 4) Assertion heuristic: Finally, we filter out any tests that do not contain any of the substrings “assert”, “verify”, or “fail”.
+These substrings are common indicators of meaningful test logic in Java-based testing frameworks.
+Tests lacking these keywords are often empty, incomplete, disabled, or rely on custom test runners and frameworks that bypass standard test assertions.
+To avoid misleading results—such as treating a test that silently does nothing as correct—we conservatively discard such test cases.
+This results in a total of 3266 runnable tests.
+We package it as Methods2Testrunnable.
+We use this as the primary testing dataset.
+We package the commit, build system, and execution status as Methods2Testmeta.
 // Focal class
@@
 } fm+fc+c+m+ft+tcTargetSource
-LISTING I: Dataset components: focal method (fm), focal class name (fc), constructor signatures (c), method signatures, fields (f), test case (t), test class name (tc).
+Listing 1: Dataset components: focal method (fm), focal class name (fc), constructor signatures (c), method signatures(m), fields (f), test case (t), test class name (tc).
 Data is assembeld according to fm+fc+c+m+f+t+tc.
@@
 The code below the red line is the target output of the unit test generation.
-IV-D2 HumanEval-X
-HumanEval is a benchmarking dataset for code generation, first introduced by Chen et al.
-[41].
+4.4.3 HumanEval-X
+HumanEval is a benchmarking dataset for code generation, first introduced by [Chen2021EvaluatingLL].
 It consists of 164 hand-written programming problems written in Python.
 For each programming problem, a function signature, docstring, body, and several unit tests are provided.
-Zheng et al.
-[42] extended the dataset to multiple languages, named HumanEval-X.
-We use the Java subset of the HumanEval-X dataset to evaluate potential adverse fine-tuning effects.
-IV-E Training
+[Zheng2023CodeGeeXAP] extended the dataset to multiple languages, named HumanEval-X.
+We use the Java subset of the HumanEval-X dataset to evaluate potential catastrophic forgetting effects of fine-tuning.
+4.5 Training
 As the training dataset Methods2Testsmall only includes method signatures for non-focal methods, the unit test generation task can be considered a sequence-to-sequence (seq2seq) problem.
-For training the decoder-only models described in Section IV-C, we, therefore, model the seq2seq task as an autoregressive problem.
+Thus, to preserve the code completion capabilities of the decoder-only models described in Section 4.3, we model the seq2seq task as an autoregressive problem.
 This is done by only calculating loss based on the target code, i.e., the unit test code.
@@
 The indicator function $\mathds{1}_{i}$ is used to ignore the source in the loss computation.
-TABLE I: Model-agnostic hyperparameters for training.
+Table 2: Model-agnostic hyperparameters for training.
 Hyperparameter Method Value
-[…]
+Common
 Optimizer - AdamW
@@
 Prompt tuning 3E-3
-[…]
+Method specific
 Alpha LoRA 32
@@
 Virtual tokens Prompt tuning 20
-We use the Transformers [43] and PEFT libraries [44] from Hugging Face to conduct all training.
-We use NVIDIA A100 GPUs and leverage DeepSpeed ZeRO-Offload [45] where necessary.
-We use the hyperparameters shown in Table I for training.
-Like other studies, e.g., [8], we select hyperparameter values that align with those specified in the original papers of the respective PEFT techniques.
-Weyssow et al.
-[8] fine-tuned their models for up to five epochs.
-Our results show that the losses of most of our fine-tuned models are stable after three epochs [^4] Available at https://huggingface.co/andstor/peft-unit-test-generation-experiments.
-Regarding the unspecified hyperparameters, we leave them to library defaults.
-Following defaults of the PEFT library, rank decomposition for LoRA is applied to the attention layers.
-The same attention layers and the default feed-forward modules (see Table II) are targeted for (IA)3.
-The selected models have context windows ranging from 2048 to 16,384.
-To make the experiments align with each other, we limit the number of tokens to the smallest window of 2048.
-TABLE II: Model-specific hyperparameters for training.
+Table 3: Model-specific hyperparameters for training.
 Hyperparameter Method Model Value
-Targeted attention modules LoRA, (IA)3 codegen-350M-multi qkv_proj
+Targeted attention modules LoRA, (IA)3 Salesforce/codegen-350M-multi qkv_proj
 Salesforce/codegen2-1B_P qkv_proj
@@
 bigcode/starcoder2-15b q_proj, v_proj
-Targeted feedforward modules (IA)3 codegen-350M-multi fc_out
+Qwen/Qwen2.5-Coder-3B q_proj,v_proj
+Qwen/Qwen2.5-Coder-7B q_proj,v_proj
+Qwen/Qwen2.5-Coder-14B q_proj,v_proj
+Targeted feedforward modules (IA)3 Salesforce/codegen-350M-multi fc_out
 Salesforce/codegen2-1B_P fc_out
@@
 bigcode/starcoder2-15b q_proj, c_proj
-IV-F Generation
-For generating, we provide the source part of the data (see LABEL:lst:data-structure) as input to the models, while the target part (original code) is used as the ground truth.
-The generated unit test code is then compared to the ground truth to calculate their similarities.
-We implement a stopping strategy based on matching braces to generate well-formed functions.
-For incomplete unit tests, we follow the protocol from [2] where we iteratively delete lines (from bottom to top), add one to two curly brackets, and check if the syntax is valid using an ANTLR v4-based [46] Java parser [47].
-We only evaluate samples that are well-formed.
-We also apply DeepSpeed-Inference [48] to facilitate model parallel (MP) to fit the larger models that would otherwise not fit in GPU memory.
-Table III shows the hyperparameter details used for generation.
-TABLE III: Hyperparameters for generation.
+Qwen/Qwen2.5-Coder-3B down_proj
+Qwen/Qwen2.5-Coder-7B down_proj
+Qwen/Qwen2.5-Coder-14B down_proj
+We use the Transformers [wolf-etal-2020-transformers] and PEFT libraries [peft] from Hugging Face to conduct all training.
+We use NVIDIA A100 GPUs and leverage DeepSpeed ZeRO-Offload [Ren2021ZeROOffloadDB] where necessary.
+We use the hyperparameters shown in Table 2 for training.
+Like other studies, e.g.
+that of [weyssow2024exploring], we select hyperparameter values that align with those specified in the original papers of the respective PEFT techniques.
+[weyssow2024exploring] fine-tuned their models for up to five epochs.
+Our results show that the losses of most of our fine-tuned models are fairly stable after three epochs [^5] Available at https://huggingface.co/datasets/andstor/peft-unit-test-generation-experiments.
+Regarding the unspecified hyperparameters, we leave them to library defaults.
+Following the defaults of the PEFT library, rank decomposition for LoRA is applied to the attention layers.
+The same attention layers and the default feed-forward modules are targeted for (IA)3.
+An exhaustive set of selected hyperparameters is provided [^6] https://huggingface.co/datasets/andstor/peft-unit-test-generation-experiments#training-hyperparameters for reproducing our results.
+The selected models have context windows ranging from 2048 to 16,384.
+To make the experiments align with each other, we limit the number of tokens to the smallest window of 2048.
+4.6 Generation
+Table 4: Hyperparameters for generation.
 Hyperparameters Value
@@
 Max length 2048
-IV-G Metrics
-IV-G1 Syntactic Correctness
-Following Yuan et al.
-[3], we measure the syntactic correctness of the generated unit tests, ensuring that they conform to the grammar of the Java programming language.
-After the fixing described in Section IV-F, we employ the same parser as described in Section IV-F.
+For generating unit tests with the Methods2Testrunnable dataset, we provide the source part of the data (exemplified in LABEL:lst:data-structure) as input to the models to generate unit test cases.
+In order to facilitate reliable execution during later evaluation stages, the source also includes the test method function definition.
+For the evaluation of catastrophic forgetting, we use the default evaluation protocol for the HumanEval-X dataset.
+This includes generating the body of a focal method, given an function signature and docstring as the source input.
+As we use a context window of 2048, we filter out any samples where the source + target overflows.
+We implement a stopping strategy based on matching braces to generate well-formed functions.
+That is, functions that have matching start and end braces.
+For incomplete unit tests, we follow the protocol from [siddiq2024using] where we iteratively delete lines (from bottom to top), add one to two curly brackets, and check if the syntax is valid using an ANTLR v4-based [antlr4] Java parser [java-antlr4].
+We only evaluate samples that are well-formed.
+We use greedy decoding without any sampling or frequency penalty.
+Table 4 shows the hyperparameter details used for generation.
+4.7 Metrics
+4.7.1 Syntactic Correctness
+Following [yuan2023manual], we measure the syntactic correctness of the generated unit tests, ensuring that they conform to the grammar of the Java programming language.
+After the fixing described in Section 4.6, we employ the same parser as described in Section 4.6.
 This involves parsing the generated tests and verifying that they are free of syntax errors.
 We then report the percentage of unit tests with valid syntax for each model.
-IV-G2 Similarity
-BLEU score [49] is a very common metric used to evaluate natural language machine translation.
+4.7.2 Similarity
+BLEU score by [Papineni2002BleuAM] is a very common metric used to evaluate natural language machine translation.
 It measures the similarity of the machine-translated text to the reference translations.
-However, Ren et al.
-[50] recognized that it neglects important syntactic and semantic features of programming languages.
-To remedy this, they developed CodeBLEU [50].
+However, [Ren2020CodeBLEUAM] recognized that the BLEU score neglects important syntactic and semantic features of programming languages.
+To remedy this, they developed CodeBLEU [Ren2020CodeBLEUAM].
 Based on n-gram matching used in BLEU, they further augment the scoring by leveraging abstract syntax trees (AST) and code semantics via data flow.
 To evaluate the primary performance of the various PEFT methods, we use CodeBLEU to measure the similarity between the generated unit test code and the ground truth.
-IV-G3 Coverage
-We also execute the generated tests for the HumanEval-Xjava, and calculate code coverage.
-Specifically, we calculate the pass@1 metric [41] and instruction and branch coverage using the Java Code Coverage Library JaCoCo [51].
-Instruction coverage measures the percentage of executed instructions, while branch coverage evaluates the percentage of executed control flow branches in the program.
-V Experimental Results
+In our experiments, we use the tokenizer_13a  [^7] https://github.com/mjpost/sacrebleu/blob/master/sacrebleu/tokenizers/tokenizer_13a.py from SacreBLEU by [bojar2018call], equivalent to the standard mteval-v13a tokenizer used by the Conference on Machine Translation (WMT).
+4.7.3 Passing Rate
+We use pass@1, a special case of the pass@k metric introduced by [Chen2021EvaluatingLL], to assess the correctness of generated outputs.
+In our setting, given a single focal method and one corresponding generated test case, pass@1 measures the probability that the test passes when executed.
+This reflects a realistic deployment scenario where there is only one chance to provide the user with a useful generated test case.
+Since our generation does not involve sampling, this also ensures reproducibility.
+4.7.4 Coverage
+To evaluate the effectiveness of the tests generated using Methods2Testrunnable, we calculate the coverage of the tests that pass.
+Specifically, we measure instruction coverage and branch coverage.
+These metrics provide a good indication of how thoroughly the test cases exercise the method under test.
+4.7.5 Mutation Score
+We evaluating the semantic strength of the generated tests by conducting mutation testing of the .
+Mutation testing evaluates the fault-detection capability of test cases by introducing small artificial defects, so-called mutants, into the program under test and checking whether the generated tests can detect them.
+A mutant is considered killed if the test fails when the mutation is present.
+We compute the mutation score as the number of mutants killed divided by the total number of mutants, following standard practice in software testing research [jia2011an].
+4.8 Test Execution
+To safely execute untrusted code, we utilize Docker containers that isolate the test environment.
+Java tests from our datasets are executed using the Apache Maven build system [apache-maven:3.8.1].
+For the Methods2Testrunnable dataset, the Maven configuration is provided by cloning the test’s repository.
+For HumanEval-X, we use a hardcoded setup of Maven, and write the test and focal method to files.
+To measure code coverage, we employ the JaCoCo Java Code Coverage Library [jacoco:0.8.3].
+For conducting mutation testing, we use the JVM mutation testing system Pitest [pitest:1.16.0].
+Pitest inserts mutants into jvm by re-writing the class after it has loaded.
+We use the following default mutators: “Conditionals Boundary”, “Increments”, “Invert Negatives”, “Mat”, “Negate Conditionals”, “Void Method Calls”, “Empty returns”, “False Returns”, “True returns”, “Null returns”, and “Primitive return”.
+Prior to execution, we replace the original implementation with the generated content.
+In the case of Methods2Testrunnable, this involves replacing the test method.
+For HumanEval-X, it involves replacing the focal method.
+We then instruct Maven to execute the single corresponding unit test method and rely on Maven to retrieve any dependencies and build required modules.
+Test execution results are obtained from the XML output produced by the standard Maven Surefire Plugin [mavensurefireplugin].
+A test run may conclude in several ways.
+It may “succeed” or “fail” based on assertion outcomes.
+It may also result in an error.
+Errors typically arise from unexpected runtime issues, such as unhandled exceptions during execution—an “interruption”.
+Beyond execution-specific issues, the environment can also lead to problems.
+These include “compilation errors”—often due to missing references, type mismatches, or unfulfilled dependencies.
+Although our datasets generally exclude unbuildable codebases (see section 4.4.2), compilation errors can still occur with generated content.
+Other environmental exceptions may arise during container setup, file I/O, or subprocess invocation, causing some tests to fail to run.
+5 Experimental Results
 This section presents the detailed results of each research question.
-V-A RQ1: Tuning method performance
-We present the performance results of the unit test generation for the various models and tuning methods in Table IV.
-TABLE IV: Comparison of syntactical validity and CodeBLEU scores from experiments using different tuning methods across various models on testing split of Methods2Testsmall and HumanEval-Xjava datasets.
-Model Method Trainable
+5.1 RQ1: Tuning method performance
+We begin by describing the performance results for each evaluation metric on the Methods2Testrunnable dataset, as shown in Section 5.1(b).
+We then explain the results to test the catastrophic forgetting effects on code generation (shown in Section 5.1(c)) using the HumanEval-X dataset.
+Evaluation metrics experiment results using different tuning methods across various models.
+(a) (b) (c)
+Method Trainable
 params
-Methods2Testsmall HumanEval-Xjava
-Valid syntax
-CodeBLEU
+Methods2Testrunnable HumanEval-Xjava
 Valid syntax
@@
 pass@1
-Instr Cov
-Branch Cov
-CodeGen-350M-multi None 0 95.43% 0.2170 100% 0.3608 0.0671 97.33% 89.77%
-Full fine-tuning 304.23M 97.87% 0.2988 100% ( 0.3293 ) ( 0.0366 ) 100% ( 83.33% )
-LoRA 1.31M ( 95.22% ) 0.2553 100% 0.3907 0.0671 98.69% ( 89.29% )
-(IA)3 0.14M 95.53% 0.2266 100% ( 0.3583 ) ( 0.0549 ) 97.69% 94.32%
-Prompt tuning 0.02M 96.03% 0.2208 100% ( 0.3290 ) ( 0.0427 ) ( 96.56% ) 91.67%
-CodeGen2-1B None 0 0% 0 0% 0 0 0% 0%
-Full fine-tuning 1,015.31M 76.73% 0.1474 5.49% 0.0359 0 0% 0%
-LoRA 2.10M 41.16% 0.0484 8.54% 0.0117 0 0% 0%
-(IA)3 0.23M 1.52% 0.2553 0% 0 0 0% 0%
-Prompt tuning 0.04M 66.63% 0.2568 7.93% 0.2547 0 0% 0%
-StarCoder2-3B None 0 85.23% 0.1543 100% 0.4264 0.3152 9.89% 85.67%
-Full fine-tuning 3,030.37M 96.71% 0.2786 100% 0.4969 0.3494 99.40% 86.37%
-LoRA 4.55M 97.11% 0.2901 100% ( 0.4169 ) 0.3675 99.19% 58.84%
-(IA)3 0.47M 87.43% 0.2513 100% ( 0.4250 ) ( 0.2744 ) 99.67% ( 83.81% )
-Prompt tuning 0.06M 86.43% 0.1742 100% 0.4309 ( 0.2470 ) 99.6% ( 75.85% )
-CodeGen2-3.7B None 0 0% 0 0% 0 0 0% 0%
-Full fine-tuning 3,641.17M 50.51% 0.1006 73.78% 0.2621 0 0% 0%
-LoRA 4.19M 52.24% 0.0997 40.24% 0.1384 0 0% 0%
-(IA)3 0.46M 0% 0 0% 0 0 0% 0%
-Prompt tuning 0.08M 23.50% 0.2562 0% 0 0 0% 0%
-CodeLlama-7B None 0 97.66% 0.3107 99.39% 0.4861 0.3293 98.33% 84.46%
-Full fine-tuning 6,607.41M ( 96.44% ) ( 0.3012 ) 100% 0.4994 0.3373 98.95% 86.37%
-LoRA 8.39M ( 97.36% ) 0.3277 99.39% ( 0.4291 ) ( 0.3129 ) 99.61% ( 72.47% )
-(IA)3 0.61M ( 97.05% ) ( 0.3011 ) 100% ( 0.4802 ) ( 0.3232 ) 98.77% 84.72%
-Prompt tuning 0.08M ( 95.93% ) ( 0.2885 ) 99.39% ( 0.4617 ) ( 0.2761 ) 98.38% ( 82.25% )
-CodeGen2-7B None 0 96.95% 0.2848 100% 0.4736 0.2256 98.31% 81.45%
-Full fine-tuning 6,862.87M 97.56% 0.3107 100% ( 0.4398 ) ( 0.1280 ) 99.75% ( 70.00% )
-LoRA 8.39M 97.87% 0.3164 100% ( 0.4636 ) ( 0.2073 ) ( 98.06% ) ( 75.35% )
-(IA)3 0.92M 97.36% 0.2904 100% 0.4898 ( 0.1829 ) 98.55% ( 79.50% )
-Prompt tuning 0.08M ( 96.64% ) ( 0.2775 ) 100% ( 0.4407 ) ( 0.2012 ) 99.10% ( 69.40% )
-StarCoder2-7B None 0 84.13% 0.1610 100% 0.4027 0.3758 99.07% 83.16%
-Full fine-tuning 7,173.92M 97.21% 0.3009 100% 0.4389 ( 0.3675 ) 99.15% 90.80%
-LoRA 7.34M 96.91% 0.3068 100% 0.5179 ( 0.3394 ) 99.35% 87.06%
-(IA)3 0.75M 94.83% 0.2903 100% 0.4213 ( 0.3697 ) 99.40% 88.39%
-Prompt tuning 0.09M ( 83.03% ) 0.3030 100% 0.5057 ( 0.3476 ) 99.38% 86.02%
-StarCoderBase None 0 84.63% 0.1563 98.78% 0.4338 0.2963 99.07% 81.48%
-Full fine-tuning 15,517.46M 96.81% 0.3123 100% 0.4830 0.3293 99.16% ( 75.20% )
-LoRA 8.03M 95.71% 0.3152 98.78% ( 0.3905 ) 0.2963 99.07% ( 73.89% )
-(IA)3 1.24M 84.63% ( 0.1553 ) 98.78% 0.4344 ( 0.1562 ) ( 98.72% ) 81.48%
-Prompt tuning 0.12M 85.73% ( 0.1518 ) ( 78.05% ) ( 0.2315 ) 0.3025 99.76% ( 67.62% )
-StarCoder2-15B None 0 85.43% 0.1898 100% 0.3724 0.4085 98.93% 87.69%
-Full fine-tuning 15,655.90M 97.90% 0.3323 ( 99.39% ) 0.4886 0.3758 99.52% 81.3%
-LoRA 12.12M 97.01% 0.3272 100% 0.4633 0.4146 98.95% 82.88%
-(IA)3 1.25M 85.43% 0.1901 100% 0.3725 0.4578 99.57% 87.89%
-Prompt tuning 0.12M 97.60% 0.3133 100% 0.5352 0.3939 99.32 % 82.89 %
-CodeGen2-16B None 0 97.87% 0.2784 100% 0.4779 0.2012 98.66% 80.56%
-Full fine-tuning 16,032.16M ( 97.56% ) 0.3383 ( 98.17% ) ( 0.3774 ) ( 0.1180 ) 99.52% ( 78.07% )
-LoRA 13.37M 98.68% 0.3186 100% ( 0.4714 ) 0.2012 98.66% 82.06%
-(IA)3 1.46M 97.87% 0.2790 100% 0.4780 0.2134 ( 97.46% ) 80.56%
-Prompt tuning 0.08M 97.97% 0.2954 100% ( 0.4679 ) 0.2195 ( 98.62% ) ( 71.07% )
-  • Bold: best-performing training method per model.
-( Parentheses ): decreased performance compared to baseline.
-Red: $<$ 50% syntactical valid samples.
-Underline: Other notable results (see in Section V-A1).
-V-A1 Syntactic Correctness
-As shown in Table IV, the percentage of syntactically correctly generated unit tests is relatively high across the board, with $>$ 80% for all except the smaller CodeGen2-1B and CodeGen2-3.7B models.
-These two cannot produce any valid unit test with the base model.
-However, after tuning, both models are able to produce up to 76.73% and 52.24% syntactically valid tests, respectively.
-The notable findings (marked with underlines in Table IV) are as follows: 1) The increase of 66.63% using prompt tuning of CodeGen2-1B, tuning only 0.02‰ (0.04 million out of 1 billion parameters) of the parameters.
-2) The increase of 52.24% using LoRA on CodeGen2-3.7B, tuning only 0.1% (4.19 million out of 3.7 billion parameters).
-Further, data in Table IV illustrate that: 1) Full fine-tuning achieves higher percentages of syntactically correct unit tests than the baseline.
-The median percentage of syntactically valid tests for the full fine-tuning is 97.01%, which is 13.69% higher than the value of the baseline (85.33%).
-2) There is a weak tendency that the more trainable parameters, the more samples are generated syntactically correctly.
-The median percentage of syntactically valid tests for the full fine-tuning is 97.01%, which is slightly higher than LoRA (96.96%), (IA)3 (91.13%), and prompt tuning (91.18%).
-However, considering that full fine-tuning often tunes parameters several thousands more times than PEFT methods, it is probably not wise to use full fine-tuning to improve the syntactic correctness of the unit tests if cost-effectiveness is a priority.
-V-A2 Similarity
-The Methods2Testsmall dataset CodeBLEU column in Table IV shows the CodeBLEU scores of the baseline, full fine-tuning, and PEFT methods.
-The notable results are: 1) LoRA archives the highest CodeBLEU scores in five of the ten models across different sizes.
-LoRA is also the only method that can positively increase the CodeBLEU score of the CodeLlama-7B model.
-2) Full fine-tuning is nearly as performant as LoRA.
-For three of the ten models, full fine-tuning scores the best.
-3) (IA)3 performs generally poorly compared to others, often resulting in very low CodeBLEU scores.
-4) Prompt tuning shows varying results, from being the worst performing on CodeLlama-7B to being the best performing on CodeGen2-1B.
-V-A3 Coverage
-From the pass@1 column in Table IV, the average mean pass@1 rate is 0.2332.
-Of the tests that pass, the code coverage is generally high, as can be seen from the two last columns in Table IV.
-The exception is where the model is not able to generate much syntactically valid code.
-This results in some coverage results being zero, e.g.
-CodeGen2-1B and CodeGen2-3.7B.
-With an average median instruction coverage of 98.81% and branch coverage of 78.55%, this shows that if the models are able to create a runnable test, they are likely to cover a significant portion of the code.
-This high level of coverage indicates that the generated tests are effective at exercising the codebase, potentially identifying edge cases, and ensuring robustness.
-The results also show that LoRA and full fine-tuning outperform other methods in several cases.
-Although (IA)3 and prompt tuning perform best in some cases, the differences between their performance and others are insignificant.
-V-A4 Resilience against catastrophic forgetting
-The results in the last five columns of Table IV show the evaluation of various PEFT tuning methods on the Java subset of HumanEval-X after they have been trained on the Methods2Testsmall dataset.
-As expected, we can see from the results that there are some cases of degradation (marked using parentheses in Table IV) when comparing the results to the baseline.
-This applies to both full fine-tuning and PEFT.
-However, the levels of degradation are, for the most part, trivial.
-We also see several cases, e.g., StarCoder2-15B, where PEFT tuning actually improves the CodeBLEU scores.
-This shows that neither tuning method is very susceptible to catastrophic forgetting, and PEFT methods might pose similarly good transfer learning capabilities to that of full-fine-tuning.
-Summary for RQ1: Full fine-tuning and PEFT methods could improve the syntactic correctness of the generated unit tests.
-PEFT methods could be better choices than full fine-tuning if cost-effectiveness to improve the syntactic correctness of the generated unit tests is a concern.
-Full fine-tuning and LoRA could significantly boost the quality of the generated unit tests measured using CodeBLU, with LoRA even surpassing full fine-tuning in several cases.
+Instr. Cov.
+Branch Cov.
+Mutation Score
+Valid syntax
+CodeBLEU
+pass@1
+[…]
+(a) (b) (c)
+Method Trainable
+params
+Methods2Testrunnable HumanEval-Xjava
+Valid syntax
+CodeBLEU
+pass@1
+Instr. Cov.
+Branch Cov.
+Mutation Score
+Valid syntax
+CodeBLEU
+pass@1
+[…]
+None 0 0.96 0.17 0.09 0.52 0.12 0.23 1.0 0.28 0.03
+Fine-tuning 356,712,448 0.98 0.22 0.14 (0.48) 0.16 0.83 1.0 (0.24) (0.02)
+LoRA 1,310,720 0.96 0.19 0.15 (0.48) 0.15 0.41 1.0 0.29 (0.02)
+(IA)3 143,360 (0.95) 0.17 0.13 (0.5) 0.13 0.45 1.0 0.28 0.03
+Prompt tuning 20,480 0.96 0.17 0.12 (0.43) 0.13 0.44 1.0 (0.25) (0.02)
+CodeGen2-1B
+None 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Fine-tuning 1,015,306,240 0.75 0.1 0.04 0.31 0.09 0.11 0.05 0.01 0.0
+LoRA 2,097,152 0.38 0.03 0.12 0.01 0.0 0.74 0.09 0.01 0.0
+(IA)3 229,376 0.02 0.26 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Prompt tuning 40,960 0.7 0.26 0.0 0.0 0.0 0.0 0.08 0.25 0.0
+StarCoder2-3B
+None 0 0.93 0.1 0.53 0.14 0.06 0.52 1.0 0.41 0.22
+Fine-tuning 3,030,371,328 0.96 0.23 (0.17) 0.62 0.22 (0.38) 1.0 0.41 (0.19)
+LoRA 4,546,560 0.98 0.24 (0.19) 0.59 0.2 0.53 1.0 (0.32) (0.16)
+(IA)3 468,480 0.94 0.21 (0.27) 0.41 0.14 (0.39) 1.0 0.42 0.24
+Prompt tuning 61,440 0.93 0.1 0.54 0.15 0.06 (0.28) 1.0 (0.38) (0.18)
+Qwen2.5-Coder-3B
+None 0 0.97 0.21 0.18 0.53 0.19 0.0 1.0 0.42 0.34
+Fine-tuning 3,085,938,688 0.98 0.25 (0.15) 0.53 (0.13) 0.0 1.0 (0.39) (0.21)
+LoRA 3,686,400 0.98 0.23 0.2 0.55 0.19 0.51 1.0 0.45 0.4
+(IA)3 479,232 0.97 0.22 0.19 0.6 0.22 0.64 1.0 0.46 0.38
+Prompt tuning 40,960 0.97 0.22 0.2 0.53 0.2 0.55 (0.99) (0.36) (0.31)
+CodeGen2-3.7B
+None 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Fine-tuning 3,641,174,016 0.41 0.07 0.05 0.33 0.15 0.0 0.74 0.26 0.0
+LoRA 4,194,304 0.42 0.07 0.12 0.09 0.06 0.58 0.4 0.04 0.0
+(IA)3 458,752 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Prompt tuning 81,920 0.22 0.26 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+CodeLlama-7B
+None 0 0.98 0.24 0.18 0.6 0.23 0.14 0.99 0.4 0.21
+Fine-tuning 6,738,546,688 0.98 0.24 0.22 0.62 (0.22) (0.0) 1.0 0.41 0.22
+LoRA 8,388,608 0.98 0.26 0.22 0.61 0.24 0.51 0.99 (0.35) (0.2)
+(IA)3 614,400 0.98 0.24 0.2 (0.59) 0.24 0.36 1.0 (0.39) 0.21
+Prompt tuning 81,920 (0.97) (0.23) (0.11) (0.59) (0.13) 0.19 0.99 (0.37) (0.2)
+CodeGen2-7B
+None 0 0.99 0.22 0.18 0.57 0.21 0.35 1.0 0.38 0.13
+Fine-tuning 6,862,858,240 (0.98) 0.23 (0.15) (0.51) (0.2) (0.27) 1.0 (0.35) (0.08)
+LoRA 8,388,608 0.99 0.24 (0.17) 0.6 0.21 0.41 1.0 (0.37) (0.12)
+(IA)3 917,504 (0.98) 0.22 0.18 0.58 0.22 0.52 1.0 0.39 0.13
+Prompt tuning 81,920 (0.98) 0.22 0.19 (0.54) 0.21 0.51 1.0 (0.36) (0.12)
+StarCoder2-7B
+None 0 0.92 0.11 0.53 0.12 0.05 0.37 1.0 0.35 0.23
+Fine-tuning 7,173,923,840 0.97 0.24 (0.2) 0.6 0.2 (0.14) 1.0 0.43 (0.22)
+LoRA 7,340,032 0.97 0.24 (0.21) 0.56 0.2 0.53 1.0 0.42 (0.22)
+(IA)3 753,664 0.95 0.24 (0.2) 0.62 0.22 0.41 1.0 0.42 (0.21)
+Prompt tuning 92,160 0.93 0.25 (0.2) 0.54 0.19 0.43 1.0 0.42 0.24
+Qwen2.5-Coder-7B
+None 0 0.97 0.23 0.2 0.59 0.21 0.49 1.0 0.39 0.37
+Fine-tuning 7,615,616,512 (0.94) (0.19) (0.14) (0.57) (0.18) (0.0) 1.0 0.41 (0.18)
+LoRA 5,046,272 0.98 0.25 0.23 0.62 0.21 0.51 1.0 (0.36) (0.32)
+(IA)3 645,120 0.98 0.24 0.24 0.64 0.21 (0.39) 1.0 0.39 0.38
+Prompt tuning 71,680 (0.96) 0.23 0.21 0.63 0.22 (0.36) 1.0 0.41 0.41
+Qwen2.5-Coder-14B
+None 0 0.98 0.25 0.24 0.58 0.2 0.4 1.0 0.37 0.3
+Fine-tuning 14,770,033,664 (0.96) (0.24) (0.2) (0.55) (0.17) (0.12) 1.0 0.43 0.35
+LoRA 12,502,912 (0.97) 0.26 0.24 0.61 0.21 0.53 1.0 0.41 0.38
+(IA)3 958,464 0.98 0.26 0.24 0.61 0.2 (0.27) 1.0 0.39 0.34
+Prompt tuning 102,400 0.98 0.25 (0.2) 0.59 0.2 0.54 1.0 0.39 0.38
+StarCoderBase
+None 0 0.92 0.11 0.48 0.15 0.06 0.28 0.99 0.4 0.2
+Fine-tuning 15,517,456,384 0.97 0.26 (0.22) 0.58 0.22 0.41 1.0 (0.39) 0.22
+LoRA 8,028,160 0.92 0.11 0.48 0.15 0.06 0.47 0.99 (0.37) 0.2
+(IA)3 1,239,040 0.92 0.11 0.48 0.16 0.06 0.46 0.99 0.4 0.2
+Prompt tuning 122,880 (0.91) 0.13 (0.34) 0.2 0.09 (0.25) (0.78) (0.22) (0.1)
+StarCoder2-15B
+None 0 0.93 0.13 0.53 0.17 0.07 0.37 1.0 0.34 0.25
+Fine-tuning 15,655,899,136 0.98 0.25 (0.19) 0.49 0.17 (0.03) (0.99) 0.4 (0.23)
+LoRA 12,124,160 0.98 0.26 (0.21) 0.61 0.22 (0.16) 1.0 0.38 0.26
+(IA)3 1,249,280 0.93 0.13 0.53 0.17 0.07 0.43 1.0 0.34 0.25
+Prompt tuning 122,880 0.98 0.26 (0.18) 0.62 0.24 0.52 1.0 0.44 0.28
+CodeGen2-16B
+None 0 0.98 0.23 0.18 0.57 0.22 0.7 1.0 0.39 0.13
+Fine-tuning 16,032,155,648 0.99 0.24 (0.13) (0.51) (0.15) (0.05) (0.98) (0.28) (0.07)
+LoRA 13,369,344 0.99 0.25 (0.17) (0.52) (0.19) (0.29) 1.0 (0.37) 0.14
+(IA)3 1,462,272 0.98 0.23 0.19 0.57 (0.21) (0.53) 1.0 0.39 0.13
+Prompt tuning 122,880 0.98 0.23 0.19 0.61 0.24 (0.54) 1.0 (0.37) 0.13
+5.1.1 Syntactic Correctness
+As shown in Section 5.1(b), the percentage of syntactically correctly generated unit tests is relatively high across the models, with >> 91% for all except the smaller CodeGen2-1B and CodeGen2-3.7B models (marked in red in Section 5.1).
+These two cannot produce any valid unit tests with the base model.
+However, after tuning, both models are able to produce up to 75% and 42% syntactically valid tests, respectively.
+Notable findings from Section 5.1(b)) are as follows: 1) The increase of 70% using prompt tuning of CodeGen2-1B, tuning only 0.02‰ (40 thousand out of 1 billion parameters) of the parameters.
+2) The increase of 42% using LoRA on CodeGen2-3.7B, tuning only 0.1% (4.19 million out of 3.7 billion parameters) of the parameters.
+Further, data in Section 5.1(b) illustrate that: 1) Full fine-tuning achieves higher percentages of syntactically correct unit tests than the baseline.
+The median value of syntactically valid tests for the full fine-tuning is 0.97, which is 1% higher than the value of the baseline (0.96).
+2) There is a slight tendency for models with more trainable parameters to generate a higher proportion of syntactically correct unit tests.
+3) The difference in syntactical validity between fine-tuning and PEFTs is not very prominent.
+The median values of syntactically valid tests of PEFT methods are: LoRA 0.98, (IA)3 0.95, and prompt tuning 0.96.
+However, considering that full fine-tuning often tunes parameters several thousand more times than PEFT methods, it is probably not wise to use full fine-tuning to improve the syntactic correctness of the unit tests if cost is a concern.
+5.1.2 Similarity
+The Methods2Testrunnable dataset CodeBLEU column in Section 5.1(b) shows the CodeBLEU scores of the baseline, full fine-tuning, and PEFT methods.
+Due to the low number of syntactically valid tests generated from the CodeGen2-1B and CodeGen2-3.7B models, we exclude these models from further analysis.
+The median CodeBLEU scores are: baseline 0.21, fine-tuning 0.24, LoRA 0.24, (IA)3 0.22, and prompt tuning 0.23.
+The notable results are: 1) LoRA achieve the highest CodeBLEU scores in seven of the eleven valid models, improving the median CodeBLEU value by 14% over the baseline.
+2) Prompt tuning is also the only PEFT method that shows slight negative results.
+5.1.3 Passing rate
+Table 5: Test execution statuses for StarCode2-7B.
+Method Succ. Failed Interrupt CompErr NoAssert
+None 53% 3% 3% 16% 26%
+Fine-tuning 20% 14% 9% 51% 6%
+LoRA 21% 15% 9% 51% 5%
+(IA)3 20% 14% 10% 49% 7%
+Prompt tuning 20% 15% 7% 52% 6%
+As explained in section 4.8, we consider a test as passing if it executes successfully without errors, as determined by a test oracle.
+Any tests that did not run due to an environmental exception (see section 4.8) are excluded from analysis because such cases did not reflect the quality of the generated test cases.
+As we use Methods2Testrunnable to generate and evaluate tests, the original focal method under test serves as a proxy for the oracle, effectively acting as the ground truth.
+While we do not assume that the methods under test are bug-free, we trust that all methods from the Methods2Test dataset are at least passable and thus suitable as oracles for evaluating generated tests.
+However, using the method as the oracle of test passing introduces a limitation: tests performing no assertion checks may still return passing results, which are misleading.
+To address this issue, we heuristically mark the generated tests as failed (even if their executions return passing results) if they do not contain any of the substrings “assert”, “verify”, or “fail”, which are indicative of meaningful assertions or validation logic in the test body.
+From the pass@1 column in Section 5.1(b), we calculated individual median pass@1 values and got the results: baseline 0.20, fine-tuning 0.17, LoRA 0.21, (IA)3 0.20, and prompt tuning 0.20.
+The baseline achieves the highest pass@1 rate on five of the eight models, with a significantly higher median score than all tuning methods.
+To understand this phenomenon, we analyzed and categorized different reasons for tests not returning a passing result.
+Table 5 shows an example of statistical analysis results of the execution status of test cases generated by the StarCode2-7B model.
+We found that the primary cause of the low number of successful test executions after fine-tuning is compilation errors.
+Although most of the generated content is syntactically valid (see section 5.1.1), compilation errors still happen, mainly due to the generated content calling nonexistent methods or having type mismatches.
+Most of these compilation errors stem from repetition or boilerplate patterns that superficially satisfy the heuristic of unit test generation without actually validating the method’s behavior.
+5.1.4 Coverage
+Among the tests returning passing results, the individual median values of the instruction coverage are: baseline 0.53, fine-tuning 0.55, LoRA 0.59, (IA)3 0.58, and prompt tuning 0.54.
+The individual median values of the branch coverage are: baseline 0.19, fine-tuning 0.18, LoRA 0.20, (IA)3 0.21, and prompt tuning 0.20.
+These numbers show that the code coverage is generally high among the tuned models.
+It means that if the tuned models can create runnable tests, the tests are likely to cover a significant portion of the code to potentially identify edge cases.
+The results also show that LoRA and full fine-tuning outperform other methods, although the differences between their performance and other PEFT methods are insignificant.
+5.1.5 Mutation score
+Of the tests that pass and returning passing results, the individual median values of the mutation scores are: baseline 0.37, fine-tuning 0.12, LoRA 0.51, (IA)3 0.43, and prompt tuning 0.44.
+A notable finding is that full fine-tuning underperforms the baseline, suggesting that extensive parameter updates may diminish the model’s ability to generate fault-revealing tests.
+Conversely, all PEFT methods outperform the baseline.
+LoRA is the most performant PEFT method in five of the eleven models, improving the median mutation scores value by 38% over the baseline.
+5.1.6 Resilience against catastrophic forgetting
+The results in Section 5.1(c) show the performance of the baseline and tuned models on a normal code generation task using the Java subset of the HumanEval-X dataset.
+The results show some cases of degradation (marked using parentheses in Section 5.1(c)) of code generation when comparing the results of the tuned models with the baseline model.
+This applies both to full fine-tuning and PEFT.
+However, the levels of degradation are, for the most part, trivial and inconsistent.
+The median CodeBLEU scores are: baseline 0.39, fine-tuning 0.40, LoRA: 0.37, (IA)3 0.39, and prompt tuning 0.37.
+The median pass@1 values are: baseline 0.22, fine-tuning 0.21, LoRA 0.20, (IA)3 0.21, and prompt tuning 0.20.
+In several cases, e.g.
+StarCoder2-15B, PEFT tuning actually improves the CodeBLEU scores and pass@1 value.
+This shows that tuning methods are mostly resilient against catastrophic forgetting, and PEFT methods exhibit transfer learning capabilities comparable to those achieved through full fine-tuning.
+Summary for RQ1: PEFT could be a better choice than full fine-tuning if cost-effectiveness to improve the syntactic correctness of the generated unit tests is a concern.
+LoRA can generate test cases as good as full fine-tuning, with consistently better mutation score.
+The tuned models may generate fewer executable test cases than baseline models, although the executable ones are better at test coverage.
 Further, PEFT methods are mostly resilient against catastrophic forgetting.
-V-B RQ2: PEFT training resource utilization effectiveness
-$10^{6}$$10^{8}$$10^{10}$rainable parametersCodeGen-350M-multiCodeGen2-1BStarCoder2-3BCodeGen2-3.7BCodeLlama-7BCodeGen2-7BStarCoder2-7BStarCoderBaseStarCoder2-15BCodeGen2-16BFull fine-tuningLoRA(IA)3Prompt tuning
-Figure 2: Comparison of the number of trainable parameters for each tuning method. Models are sorted by size.
+5.2 RQ2: PEFT training resource utilization effectiveness
 One of the main advantages of using PEFT methods is that they require significantly less computing during training, mainly because of reduced backpropagation calculations.
 The fewer trainable parameters, the more efficient.
-Results in Figure 2 show a significant reduction of trainable parameters of PEFT methods compared to full fine-tuning.
+As seen from the second column in Section 5.1(a), PEFT methods show a significant reduction of trainable parameters compared to full fine-tuning.
 This is also essential to fit big models into GPU memory.
-For example, we could train the biggest 16 billion parameter model using a single NVIDIA A100 80GB GPU for LoRA, whereas we needed 6 x NVIDIA A100 40GB GPU for full fine-tuning.
-CodeGen-350M-multiCodeGen2-1BStarCoder2-3BCodeGen2-3.7BCodeLlama-7BCodeGen2-7BStarCoder2-7BStarCoderBaseStarCoder2-15BCodeGen2-16B1.01.0-1.0- 1.00.50.5-0.5- 0.50.00.00.00.00.50.50.50.51.01.01.01.0Relative importanceLoRA(IA)3Prompt tuning
-Figure 3: Stacked bar chart of symmetrically normalized CodeBLEU score increase and decrease, relative to the fraction of trained parameters for each model.
-Positive values mean an increased CodeBLEU score, while negative values indicate a decreased CodeBLEU score.
-Figure 3 shows the CodeBLEU score of each tuning method relative to the fraction of trainable parameters, symmetrically normalized between -1 and 1.
-As shown in Figure 3, prompt tuning has the highest efficiency among seven (i.e., CodeGen-350M-multi, CodeGen2-1B, StarCoder2-3B, CodeGen2-3.7B, StarCoder2-7B, StarCoder2-15B, and CodeGen2-16B) of the ten models and is the method that you get the ”most bang for your buck.” Moreover, prompt tuning has the most potential effect on models with large sizes, i.e., StarCoder2-15B and CodeGen2-16B.
-This can be explained by the increased capacity of larger models to understand better and leverage the additional context provided by prompt tuning.
+For example, we could train the largest 16 billion parameter model using a single NVIDIA A100 80GB GPU for LoRA, whereas we needed 6 x NVIDIA A100 40GB GPUs for full fine-tuning.
+Figure 2: Stacked bar chart of symmetrically normalized metrics value differences (increase or decrease) between baseline and tuned models, divided by the percentage of trainable parameters of models, on the Methods2Testrunnable dataset.
+Figure 2 shows the results of the metrics (valid syntax, CodeBLEU, pass@1, instruction coverage, branch coverage, and mutation score) for each tuning method relative to the percentage of trainable parameters, symmetrically normalized between -1 and 1.
+As shown in Figure 2, prompt tuning has the highest efficiency for every metric and is the method that you get the “most bang for your buck.” Moreover, prompt tuning has the most potential cost-effectiveness on models with large sizes, i.e., Qwen2.5-Coder-14B, StarCoderBase, StarCoder2-15B, and CodeGen2-16B.
+This can be explained by the increased capacity of larger models to better understand and leverage the additional context provided by prompt tuning.
 However, prompt tuning does not show consistent superiority for all the models.
-The method contributes to negative performance in three models (CodeLlama-7B, CodeGen2-7B, and StarCoderBase).
-For the three smallest models (CodeGen-350M-multi, CodeGen2-1B, and StarCoder2-3B), (IA)3 clocks in at second place, which provides better results than LoRA.
-However, (IA)3 fails to improve the three largest models.
+The method contributes to negative performance in three models (StarCoder2-3B, CodeLlama-7B, and CodeGen2-7B).
+For the four smallest models (Qwen2.5-Coder-3B, CodeGen-350M-multi, CodeGen2-1B, and StarCoder2-3B), (IA)3 clocks in at second place, which provides better results than LoRA.
+However, (IA)3 generally fails to efficiently improve the three largest models.
 LoRA achieves a relatively low efficiency compared to the other PEFT methods.
-However, it does not lead to a negative impact on the performance of all models.
+However, with the exception of pass@1, it does not lead to any significant negative impact on the performance of any of the models.
 Full fine-tuning obviously has the worst efficiency, as it needs to train all the model parameters.
-Summary for RQ2: PEFT methods offer significantly reduced computational requirements compared to full fine-tuning and are ranked by resource utilization effectiveness as follows: prompt tuning $>$ (IA)3 $>$ LoRA $>$ full fine-tuning.
+Summary for RQ2: PEFT methods offer significantly reduced computational requirements compared to full fine-tuning and are ranked by resource utilization effectiveness as follows: prompt tuning >> (IA)3 >> LoRA >> full fine-tuning.
 Prompt tuning demonstrates the highest returns on investment for most models but also carries risks of negative performance impact.
-VI Discussion
+6 Discussion
 This section highlights our novel insights by comparing our results with related work. We also discuss possible threats to the validity of our study and how we mitigate them.
-VI-A Comparison with related works
-Unlike existing studies, we investigated using PEFT and LLM to generate complete unit tests because the effectiveness of PEFT methods is task-dependent.
-Our results provide novel insights into their performance in the context of unit test generation.
-For example, Ayupov and Chirkova [5] conclude that LoRA is less effective than full fine-tuning for code generation.
-We show contradictory results that LoRA is more effective than fine-tuning in unit test generation.
-Weyssow et al.
-[8] explored generating code from natural language instructions and concluded that LoRA is the most effective one in the PEFT methods they evaluate.
-Our results show that prompt tuning is the most effective in terms of cost and resource utilization in unit test generation, although LoRA is more effective without considering the cost aspects.
-Compared to the existing studies, we also performed a more comprehensive evaluation, as we invested significant efforts to compare all PEFT methods to full fine-tuning across all model sizes (up to 16B parameters).
-In contrast, Ding et al.
-[24] only evaluated PEFT using the 355 million parameter RoBERTaLARGE model, while Weyssow et al.
-[8] only did full fine-tuning for smaller models up to maximum 350 million parameters.
-Several studies have leveraged LLMs for generating unit tests [1, 2, 3, 4, 29].
+6.1 Comparison with related studies
+As the effectiveness of PEFT methods is task-dependent, our results provide novel insights into their performance in the context of unit test generation.
+[Ayupov2022ParameterEfficientFO] conclude that LoRA is less effective than full fine-tuning for code generation.
+We show contradictory results that LoRA could be as good as or even more effective than fine-tuning in unit test generation.
+[weyssow2024exploring] explored generating code from natural language instructions and concluded that LoRA is the most effective among the PEFT methods they evaluated.
+Our results show that prompt tuning is the most effective in terms of cost and resource utilization in unit test generation, although LoRA provides better performance without considering the cost aspects.
+Compared to the existing studies, we also performed a more comprehensive evaluation, as we invested significant efforts to compare all PEFT methods to full fine-tuning across all model sizes (up to 16 billion parameters).
+In contrast, [Ding2023ParameterefficientFO] only evaluated PEFT using the 355 million parameter RoBERTaLARGE model, while [weyssow2024exploring] only did full fine-tuning for smaller models up to a maximum of 350 million parameters.
+While existing works on code generation or test generation have demonstrated promising results, they are often limited to artificially constructed benchmarks such as HumanEval.
+These datasets are typically small, synthetically curated, and lack the complexity of real-world software systems.
+As our results show, such controlled settings also overlook critical challenges that arise in real-world scenarios, such as missing dependencies, inconsistent, or erroneous code.
+Defects4J [just2014defects4j], one of the more realistic benchmarks available, contains only 17 repositories.
+This starkly contrasts the scale and messiness of actual industrial or open-source projects.
+As a result, prior work such as [alagarsamy2024a3test] may significantly overestimate the robustness and generalizability of models when deployed beyond synthetic testbeds.
+Our evaluation, grounded in more realistic settings using Methods2Test [tufano2021unit], highlights these limitations and emphasizes the need for evaluation frameworks that reflect the true complexity of software development.
+Several studies have leveraged LLMs for generating unit tests [tufano2021unit, siddiq2024using, yuan2023manual, schäfer2023empirical, steenhoek2023reinforcement].
 However, none of the works investigate the potential performance of using PEFT compared to full fine-tuning.
-Tufano et al.
-[1] automatically generated unit tests and reported 95% syntactic correctness when training BART large (400 million parameters) [52] on Methods2Test.
-With PEFT methods, we can generate an average median of 94.03% syntactically correct unit tests.
-Training between 0.02 million and 13 million parameters on 1% of the same data showcasing the potential powerful performance of PEFT methods.
-Our results also illustrate that LoRA surpasses full fine-tuning in terms of boosting the quality of the generated unit tests measured using CodeBLU in several cases.
-VI-B Implications to Academia and Industry
+[tufano2021unit] automatically generated unit tests and reported 95% syntactic correctness when training BART large (400 million parameters) [lewis2019bart] on Methods2Test.
+With PEFT methods, we can generate an average median of 96% syntactically correct unit tests.
+Training between 20 thousand and 13 million parameters on 1% of the same data, showcasing the potentially powerful performance of PEFT methods.
+Different from studies [siddiq2024using, yuan2023manual, schäfer2023empirical] using large instruction-tuned enterprise LLMs such as ChatGPT to generate test cases, this study focused on generating test cases from autoregressive base code models.
+Compared to e.g.
+[schäfer2023empirical], our best model produced 13% less statement coverage and 55% lower branch coverage.
+Achieving state-of-the-art performance was not the focus of our work.
+Instead, our work provides an empirical evaluation of PEFT methods in a more practical scenario, where organizations can efficiently customize models for their own testing needs.
+One reason for the reduced coverage is probably due to the size differences of the models.
+The largest model we evaluated had 16 billion paramaters, opposed to the lager commercial GPT-3.5-Turbo.
+Our results shown in Section 5.1 illustrate that larger models generally generate better unit tests than smaller ones.
+Another reason is probabaly due to the differences between the benchmark datasets used by our study.
+We use a real-world evaluation dataset that is more diverse, spanning 119 repositories, versus the 25 npm packages used by [siddiq2024using].
+Our analyses on the reasons for low pass@1 rates show that it is necessary to consider potential compilation and code dependency issues when using LLMs to generate executable unit tests, providing valuable novel insights to improve unit test generation using PEFT or ChatGPT.
+Opposed to related works like [siddiq2024using, yuan2023manual, schäfer2023empirical], we conduct mutation testing.
+As our results depict in section 5.1.5, measuring mutation score is crucial to ensure generated tests are actually able to reveal faults.
+6.2 Implications to Academia and Industry
 Writing unit tests is a hard and often time-consuming task [7102609].
-However, they are essential to ensure code quality, stability, and maintainability [54].
+However, they are essential to ensure code quality, stability, and maintainability [Gonzalez2017ALS].
 Automating unit test generation using LLMs can significantly reduce the time and effort required for this process, making it an attractive solution for academia and industry.
-However, as our experiments show, full fine-tuning is sometimes required to achieve adequate quality of the generated unit tests, which is computationally demanding.
+As our experiments show, full fine-tuning is sometimes required to achieve adequate quality of the generated unit tests, which is computationally demanding.
 Many LLM users may not have the necessary computing resources to fine-tune LLMs fully.
-Our results show that various PEFT methods can achieve similar performance as full fine-tuning by training a minuscule parameter fraction.
+Our results show that various PEFT methods can achieve similar performance to full fine-tuning by training a minuscule parameter fraction.
 Furthermore, our results suggest starting with prompt tuning for cost-effectiveness.
-Then, if stability issues arise, LoRA should be considered as a reliable alternative because it does not lead to negative performance impact.
-VI-C Threaths to Validity
-VI-C1 External validity
+Then, if stability issues arise, LoRA should be considered as a reliable alternative because it does not lead to a negative performance impact.
+6.3 Threats to Validity
+6.3.1 External validity
 One of the threats to external validity is the selection of models.
 We minimize this threat by selecting a broad range of models varying in size, pre-training data, and learning objectives.
-Another potential concern is the validity of the Methods2Testsmall benchmarking dataset.
-We opted for this dataset over other alternatives because it comprises real-world unit tests.
-We also validate the Java-translated version of the popularly used HumanEval code benchmarking dataset to mitigate the threat.
+As we have have spent upward of 700 GPU hours for model tuning alone, exploring an even broader range of alternative models is constrained due to cost.
+Another potential concern is the validity of the benchmarking dataset.
+We chose to use Methods2Test, which is the largest publicly available corpus of real-world Java unit test cases and corresponding focal methods.
+We believe this provides a representative evaluation setting compared to existing studies like [alagarsamy2024a3test], although future work should expand the analysis to other programming languages to ensure generalizability.
 The last threat is related to our evaluation metric.
@@
 Although CodeBLEU may not be the optimal metric to evaluate the functional correctness of unit tests completely, it provides a good tradeoff between evaluation effort and validity.
-To supplement our evaluation, we also calculate pass@1 and code coverage for the Methods2Testsmall dataset, as this contains isolated code samples that are easier to test.
-VI-C2 Internal validity
+To supplement our evaluation, we have also spent significant effort developing evaluation pipelines, as well as over 200 CPU hours for executing the test in the Methods2Testrunnable dataset, calculating pass@1, code coverage, and mutation scores.
+6.3.2 Internal validity
 The main threat to internal validity is comprised of the selection of hyperparameters.
-We use the hyperparameters reported in the introductory paper of each PEFT method.
+We use the hyperparameters reported in the introductory paper of each PEFT method to avoid introducing unnecessary biases.
 Additionally, we use greedy decoding to support reproducibility and reduce variability in the generated results.
-VII Conclusion and future work
+7 Conclusion and Future Work
 Although PEFT has been investigated to boost the quality and efficiency of code-related tasks using LLMs, they have not been studied in the context of unit test generation.
 We performed the first empirical study to compare different PEFT methods and full fine-tuning to generate unit tests.
-Our comparison involved ten LLMs of different sizes and architectures.
+Our comparison involved thirteen LLMs of different sizes and architectures.
 Our results offer valuable insights into the advantages and challenges of using PEFT and full fine-tuning for unit test generation.
 With the knowledge we’ve gathered, individuals can choose the most suitable PEFT methods or full fine-tuning based on the size and characteristics of the LLMs.
+In future work, we plan to extend the analysis perform an empirical comparison of methods for tuning instruction-based LLMs, covering both PEFT methods per se and variations in prompt templates
 In addition to unit tests, manually coding other automated test cases, such as integration and system tests, can be cost-intensive.
 In the future, we intend to explore the utilization of LLMs in conjunction with cost-effective tuning methods to automatically generate other types of tests.
-VIII Data availability
-The Methods2Testsmall dataset is available at:
-https://huggingface.co/datasets/andstor/methods2test_small
-All fine-tuned models are also made available at:
-https://huggingface.co/andstor/peft-unit-test-generation-experiments
-Experiment code and results are available at:
-https://github.com/andstor/peft-unit-test-generation-replication-package
+Acknowledgements
+The empirical results were supported through the computational resources of HPC facilities at the Norwegian University of Science and Technology (NTNU) [sjalander2019epic].
+Data Availability
+The Methods2Testsmall dataset is available at: https://huggingface.co/datasets/andstor/methods2test_small
+The Methods2Testrunnable dataset is available at: https://huggingface.co/datasets/andstor/methods2test_runnable
+The Methods2Testmeta dataset is available at: https://huggingface.co/datasets/andstor/methods2test_meta
+Metadata and links to the trained models can be found at: https://huggingface.co/datasets/andstor/peft-unit-test-generation-experiments
+Experiment code and results are available at: https://github.com/andstor/peft-unit-test-generation-replication-package